### v1

# HERA: High-efficiency Matrix Compression via Element Replacement

Abstract
Large Language Models (LLMs) have significantly advanced natural language processing tasks such as machine translation, text generation, and sentiment analysis. However, their large size, often consisting of billions of parameters, poses challenges for storage, computation, and deployment, particularly in resource-constrained environments like mobile devices and edge computing platforms. Additionally, the key-value (k-v) cache used to speed up query processing requires substantial memory and storage, exacerbating these challenges. Vector databases have emerged as a crucial technology to efficiently manage and retrieve the high-dimensional vectors produced by LLMs, facilitating faster data access and reducing computational demands.
Effective compression and quantization techniques are essential to address these challenges, as they reduce the memory footprint and computational requirements without significantly compromising performance. Traditional methods that uniformly map parameters to compressed spaces often fail to account for the uneven distribution of parameters, leading to considerable accuracy loss. Therefore, innovative approaches are needed to achieve better compression ratios while preserving model performance.
In this work, we propose HERA, a novel algorithm that employs heuristic Element Replacement for compressing matrix. HERA systematically replaces elements within the model using heuristic methods, which simplifies the structure of the model and makes subsequent compression more effective. By hierarchically segmenting, compressing, and reorganizing the matrix dataset, our method can effectively reduce the quantization error to 12.3% of the original at the same compression ratio.
1 Introduction
Large Language Models (LLMs) have revolutionized the field of natural language processing (NLP), enabling significant advancements in tasks such as machine translation, text generation, and sentiment analysis. These models, characterized by their large-scale neural network architectures and vast training datasets, have shown remarkable capabilities in understanding and generating human language. The advent of LLMs, such as OpenAI’s GPT-3 and BERT by Google, has pushed the boundaries of what machines can achieve in linguistic tasks, providing near-human performance in various applications  Brown et al. (2020); Devlin et al. (2018).
The development of LLMs is rooted in the transformer architecture, which employs self-attention mechanisms to process and produce language with high contextual relevance. This architecture has replaced previous recurrent neural network (RNN) and long short-term memory (LSTM) models due to its efficiency and ability to handle long-range dependencies in text Vaswani et al. (2017). As a result, LLMs have become the cornerstone of modern NLP, driving innovations in areas such as automated customer service, content creation, and real-time language translation.
Furthermore, the development and deployment of LLMs face significant challenges, particularly in optimizing vector database efficiency, LLM weight quantization, and key-value (k-v) cache optimization. First, as the size and complexity of LLMs increase, efficiently managing and retrieving vectors in large databases becomes critical. Optimizing vector database efficiency involves improving indexing and search algorithms to handle high-dimensional data, which is essential for tasks like semantic search and recommendation systems Johnson et al. (2019). Second, LLM weight quantization, which involves reducing the precision of model weights, is a crucial technique for making these models more storage and computation efficient without significantly degrading their performance. Quantization can drastically reduce the memory footprint and computational requirements, enabling the deployment of LLMs on resource-constrained devices. However, achieving optimal quantization while maintaining model accuracy remains a complex challenge that requires sophisticated algorithms and techniques Dettmers et al. (2021). Lastly, optimizing k-v cache efficiency is vital for speeding up inference times in LLMs. The k-v cache stores intermediate activations during the forward pass, which can be reused to avoid redundant computations. Efficient management and compression of these caches are essential to reduce latency and improve the throughput of LLMs, especially in real-time applications like chatbots and virtual assistants.
Figure 1: The distribution of LLM parameter.
To address the challenges of storage and computational efficiency in large language models (LLMs), it is crucial to develop techniques that optimize vector database efficiency, weight quantization, and key-value (k-v) cache optimization. Therefore, finding quantization algorithms that can efficiently compress and quickly decompress matrix data is essential.
In practice, when compressing parameters, the distribution of parameters is often uneven. First, the distribution is not uniform, and second, the magnitudes of different data points vary significantly. For example, the distribution of Transformer neural network weights is shown in Figure  1. In Figure  1, different rows represent the 0th layer, 8th layer, 16th layer, and 24th layer. Different columns represent the attention layer’s q, k, v, and o layers, as well as the fully connected layer’s gate, up, and down layers. The distributions of different layers are not uniformly analyzed, and the distributions vary across different layers. Previous algorithms uniformly and linearly map parameters to another compressed space. Although this method is simple, it may not be effective and can result in significant accuracy loss because it does not take into account the original data distribution. It maps both densely populated and sparse regions to the quantized space in the same way. Besides, A lines of works have been proposed to optimize vector database efficiency Ge et al. (2013); Jegou et al. (2010); Kalantidis, Avrithis (2014), quantize the model weights Dettmers et al. (2022); Polino et al. (2018); Chmiel et al. (2020); Fan et al. (2020); Zafrir et al. (2019); Wu et al. (2022); LeCun et al. (1989) and optimizate kv cache Hooper et al. (2024); Cho et al. (2024). They also does not take into account the original data distribution. Therefore, a more reasonable approach is to quantize the model based on the distribution and magnitude of the parameters, stratifying and computing accordingly.
In this work, we propose a novel algorithm named HERA to enhance the compression of large language models. The core idea of HERA is to group the relevant matrices and compress each group using clustering methods. Furthermore, HERA reorders the original dataset based on the distribution and magnitude of the parameters. This reordering is performed in multiple layers, with each layer featuring a new arrangement that optimizes the dataset’s distribution for compression. By leveraging these techniques, HERA effectively reduces data error and minimizes the discrepancy between estimated data and actual data.
We implemented a prototype system using Python. Experimental results show that, by hierarchically segmenting, compressing, and reorganizing the matrix dataset, our method can effectively reduce the quantization error to 12.3% of the original at the same compression ratio.
2 HERA
2.1 Basic algorithm
We borrow the idea from product quantization algorithm and k-means algorithm to compress the model parameter. The algorithm is divided into four steps, the first three steps is quantization process as shown in Figure  2, the last two steps is the dequantization process as show in Figure  3. The detailed process is described below.
Algorithm 1 Product Quantization Algorithm
1:$X\in\mathbb{R}^{N\times D}$triangleright Input matrix
2:Quantization Process:
3:Step 1: Space Partitioning
4:for each vector $x_{i}$ in $X$ do
5:     Divide $x_{i}$ into $M$ low-dimensional subspaces $x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{M}$
6:end for
7:Step 2: Codebook Calculation
8:for each subspace $j$ from 1 to $M$ do
9:     Apply k-means algorithm to subspace $j$ to get $k$ cluster centers
10:     Create codebook $C_{j}$ for subspace $j$ with cluster centers
11:end for
12:Step 3: Quantization
13:for each vector $x_{i}$ in $X$ do
14:     for each subspace $j$ from 1 to $M$ do
15:         Map $x_{i}^{j}$ to the nearest cluster center in $C_{j}$
16:     end for
17:end for
18:Dequantization Process:
19:Step 4: Codebook Reconstruction
20:Construct a global codebook $C$ that contains all possible combinations of cluster centers from $C_{1},C_{2},\ldots,C_{M}$
21:Step 5: Dequantization
22:for each quantized vector $q_{i}$ in the quantized dataset do
23:     Restore $q_{i}$ to its original high-dimensional space using the global codebook $C$
24:end for
  1. 1.
    Space Partitioning: Divide the high-dimensional space into a series of low-dimensional subspaces.
  2. 2.
    Codebook calculation: Use k-means algorithm to calculate codebook of low-dimensional subspaces.
  3. 3.
    Quantization: Perform independent quantization within each low-dimensional subspace, mapping data points to the nearest cluster center.
  4. 4.
    Codebook Reconstruction: Reconstruct the data using a codebook that contains all possible cluster centers.
  5. 5.
    Dequantization: Restore the vectors in the original high-dimensional space based on the vectors in the quantized low-dimensional subspaces.
Figure 2: The Quantization Process of HERA.
Figure 3: The Dequantization Process of HERA.
In data matrices where the distribution of data is uneven across different columns, directly applying the k-means algorithm for clustering may yield suboptimal results. This unevenness can distort the clustering process, as k-means assumes that all features contribute equally to the distance calculations. The differing distributions across columns mean that some features may disproportionately influence the clustering outcome, leading to biased clusters that do not accurately reflect the true structure of the data.
To mitigate these effects, it is crucial to preprocess the data appropriately. One common approach is to normalize the data so that each column contributes equally to the distance metric used by k-means. Standard normalization techniques, such as z-score normalization or min-max scaling, can be employed to transform the data into a more uniform distribution. By ensuring that all features are on a comparable scale, the k-means algorithm can perform more effectively, producing clusters that better represent the underlying patterns in the data.
Algorithm 2 Product Quantization with HERA heuristics
1:Input matrix: $X\in\mathbb{R}^{N\times D}$, number of iteration in HERA: $iter$.
2:Quantized feature matrix and feature map
3:for each element $x_{i}\in X$ do
4:     Rearrange $X$ into pairs: $X_{\text{pairs}}\in\mathbb{R}^{N\times D/2}$
5:     for each dimension $j$ do
6:         $x_{pair_{i,j}}=(x_{2i,j},x_{2i+1,j})\in X_{\text{pairs}}$
7:     end for
8:end for
9:for each pair $x_{pair_{i,j}}$ in $X_{\text{pairs}}$ do
10:     if $x_{pair_{i,j}}[1]<x_{pair_{i,j}}[2]$ then
11:         $X[iter]_{small_{i,j}}\leftarrow x_{pair_{i,j}}[1]$
12:         $X[iter]_{big_{i,j}}\leftarrow x_{pair_{i,j}}[2]$
13:         $fm[iter]_{i,j}\leftarrow 1$
14:     else
15:         $X[iter]_{small_{i,j}}\leftarrow x_{pair_{i,j}}[2]$
16:         $X[iter]_{big_{i,j}}\leftarrow x_{pair_{i,j}}[1]$
17:         $fm[iter]_{i,j}\leftarrow 0$
18:     end if
19:end for
20:for $X_{small}$ and $X_{large}$ do
21:     code[$iter$], codebook [$iter$] $\leftarrow$ PQ($X[iter]_{small}$)
22:     code[$iter$], codebook[$iter$] $\leftarrow$ PQ($X[iter]_{large}$)
23:end for
24:$iter$leftarrow $iter$ 1
25:if $iter$neqthen
26:     Goto line 1
27:end if
However, these transformations require specific parameters for each column and involve complex calculations, which can be computationally intensive and demand significant storage resources. To address this, we employ a reordering method that permutations the elements of the matrix based on their relative sizes. This approach achieves a more uniform distribution, enabling more effective application of compression operations.
Algorithm 3 Dequantization of Product Quantization with HERA heuristics
1:Quantized feature matrices: $code_{small}$, $code_{big}$, codebooks: $codebook_{small}$, $codebook_{big}$, feature maps: $fm$, number of iterations in HERA: $iter$
2:Reconstructed matrix: $\hat{X}\in\mathbb{R}^{N\times D}$
3:Initialize $\hat{X}_{pairs}\in\mathbb{R}^{N\times D/2}$
4:for each iteration from $1$ to $iter$ do
5:     for each pair $x_{pair_{i,j}}$ in $\hat{X}_{pairs}$ do
6:         $\hat{X}_{small_{i,j}}\leftarrow\text{decode}(code_{small}[iter],codebook_{small}[iter])$
7:         $\hat{X}_{big_{i,j}}\leftarrow\text{decode}(code_{big}[iter],codebook_{big}[iter])$
8:     end for
9:     for each pair $x_{pair_{i,j}}$ in $\hat{X}_{pairs}$ do
10:         if $fm[iter]_{i,j}=1$ then
11:              $x_{pair_{i,j}}[1]\leftarrow\hat{X}_{small_{i,j}}$
12:              $x_{pair_{i,j}}[2]\leftarrow\hat{X}_{big_{i,j}}$
13:         else
14:              $x_{pair_{i,j}}[2]\leftarrow\hat{X}_{small_{i,j}}$
15:              $x_{pair_{i,j}}[1]\leftarrow\hat{X}_{big_{i,j}}$
16:         end if
17:     end for
18:end for
19:for each dimension $j$ do
20:     for each element $i$ in $\hat{X}_{pairs}$ do
21:         $\hat{x}_{2i,j}\leftarrow x_{pair_{i,j}}[1]$
22:         $\hat{x}_{2i+1,j}\leftarrow x_{pair_{i,j}}[2]$
23:     end for
24:end for
The purpose of permutation is to explore all possible reordering methods within the search space. For an $N\times D$ matrix, there are $(N!)^{D}$ possible permutations, creating a vast search space.
To manage this, we employ heuristic algorithms to reduce the search space and optimize the algorithm’s complexity. Heuristic algorithms provide approximate solutions by focusing on the most promising areas of the search space, rather than exhaustively evaluating all possible permutations. These methods significantly decrease computational requirements while still delivering high-quality solutions.
By adopting this heuristic reordering strategy, we aim to achieve a more balanced distribution of data across the matrix. This not only simplifies the preprocessing steps but also enhances the efficiency of subsequent clustering and compression processes. Ultimately, this method helps overcome the challenges posed by uneven data distributions, leading to more accurate and resource-efficient clustering outcomes.
2.2 HERA heuristics
To improve quantization and avoid uneven distribution, HERA first perceives the distribution of data sizes, then rearranges the elements according to the distribution, and finally quantizes the rearranged matrix.
2.2.1 HERA Design
We observe that compressing matrix elements of similar sizes as a single column (group) yields better results.HERA uses a low-cost method to locally adjust element arrangement based on data sizes. We first introude the HERA design. The design of HERA heuristics is shown in Figure  4.
Figure 4: HERA heuristics.
HERA groups every two elements in the matrix together, rearranges the matrix elements based on the size of adjacent elements, and then quantizes the rearranged results.
Quantization: The quantization process is illustrated in the algorithm 2, the quantization process consists of multiple steps, which involve rearranging the input data, comparing elements, and quantizing both smaller and larger value sets.
The algorithm begins by rearranging the input matrix $X\in\mathbb{R}^{N\times D}$ into pairs (lines 1 in Algorithm 2). Each element $x_{i}\in X$ is grouped into $X_{\text{pairs}}\in\mathbb{R}^{N\times D/2}$ by combining every two consecutive elements along each dimension. For each dimension $j$, the pairs are formed as $(x_{2i,j},x_{2i+1,j})$ (lines 2–4 in Algorithm 2).
Next, each pair $x_{pair_{i,j}}$ is compared to separate the smaller and larger values. If the first element of the pair is smaller, it is assigned to $X_{small}$ and the second to $X_{big}$, with a feature map $fm$ recording the comparison result as 1 (lines 5–8 in Algorithm 2). Otherwise, the assignment is reversed, and the feature map records a 0 (lines 9–12 in Algorithm 2).
The quantization process proceeds by applying the PQ algorithm to both $X_{small}$ and $X_{big}$. This process is iteratively repeated for a predefined number of iterations ($iter$), refining the quantization results with each iteration (lines 13–16 in Algorithm 2).
Dequantization:The Decompression process consists of multiple steps, namely decoding the quantized values, reconstructing the pairs, and restoring the original matrix dimensions.
The algorithm begins by initializing the reconstructed matrix $\hat{X}_{pairs}\in\mathbb{R}^{N\times D/2}$ (line 1 in Algorithm 3). For each iteration, the algorithm decodes the quantized values from the codebooks for both small and large pairs (lines 2–4 in Algorithm 3). The decoded values are stored in $\hat{X}_{small}$ and $\hat{X}_{big}$ respectively.
Next, for each pair in $\hat{X}_{pairs}$, the algorithm reconstructs the pairs based on the feature map $fm[iter]$. If the feature map value is 1, the first element of the pair is assigned the decoded small value and the second element the decoded big value (lines 5–8 in Algorithm 3). If the feature map value is 0, the assignment is reversed (lines 9–12 in Algorithm 3).
Finally, the pairs are expanded back into the original dimensions to form the reconstructed matrix $\hat{X}$ (lines 13–17 in Algorithm 3).
3 Experiments
3.1 Experiment setup
Dataset: The datasets used in our experiments were generated with matrices of size $N\times D$, where $N$ is the number of samples and $D$ is the dimensionality of each sample. We employed one types of normal distribution to create these datasets.
Normal Distribution Dataset: For the normal distribution dataset, each element of the matrix was drawn from a truncated normal distribution with a mean of 0.5 and a standard deviation of 0.16. The distribution was truncated to the open interval $(0,1)$ to ensure all values remain within this range.
$\mathbf{X}_{ij}\sim\mathcal{TN}(0.5,0.16,0,1),\quad\forall i\in\{1,\ldots,N\},\quad\forall j\in\{1,\ldots,D\}$
Here, $\mathcal{TN}(\mu,\sigma,a,b)$ denotes the truncated normal distribution with mean $\mu$, standard deviation $\sigma$, truncated to the interval $(a,b)$.
Platform and implementation: We conducted our algorithm evaluations on a high-performance server equipped with an Intel Core i9-10980XE processor, featuring 18 cores and 36 threads, operating at a base frequency of 3.00 GHz. The server also includes 128GB of 3200MHz DDR4 memory and a 24.8MB L3 cache, providing robust computational capabilities. All algorithms were implemented in Python, using version 3.8.10. For each case, the experiment was repeated 100 times. In each repetition, matrices of the same size and from the same distribution were generated using different seeds.
Comparison of Algorithms and Parameter Selection
In this section, we describe the parameter selection process for our HERA algorithm and compare its performance with the Optimized Product Quantization (OPQ) algorithm.
Our HERA algorithm involves a critical parameter, the number of centroids $K_{s}$. The selection of $K_{s}$ is governed by the following memory constraint:
$K_{s}\times D\times 32+N\times M\times[(K_{s}-1)\cdot\text{bitlength}]\leq\text{memory in bits}$
where:
  • $K_{s}$ is the number of centroids,
  • $D$ is the dimensionality of each sample,
  • $N$ is the number of samples,
  • $M$ is the number of subspaces,
  • bitlength is the bit length used for encoding.
This constraint ensures that the total memory usage remains within the available memory limits. By carefully selecting $K_{s}$, we aim to optimize the performance of the HERA algorithm while adhering to memory constraints.
The Optimized Product Quantization (OPQ) algorithm was used as a benchmark for comparison. The OPQ algorithm was implemented with default parameters, providing a consistent basis for performance evaluation. The OPQ algorithm optimizes product quantization by adjusting the space partitioning and centroid assignments to minimize quantization error.
Comparison Metrics
We use MAE (Mean Absolute Error), MRE (Mean Relative Error), and MSE (Mean Squared Error) for experimental evaluation. Let $X_{(i,j)}$ and $X_{(i,j)}^{\prime}$ denote the values before and after dequantization, respectively. The following metrics are used for comparison:
$\text{MAE}=\frac{1}{N\cdot D}\sum_{i,j}\left|X_{(i,j)}-X_{(i,j)}^{\prime}\right|$
$\text{MRE}=\frac{1}{N\cdot D}\sum_{i,j}\frac{\left|X_{(i,j)}-X_{(i,j)}^{\prime}\right|}{X_{(i,j)}}$
$\text{MSE}=\frac{1}{N\cdot D}\sum_{i,j}\left(X_{(i,j)}-X_{(i,j)}^{\prime}\right)^{2}$
We performed an optimization using 1-4 iterations, corresponding to our algorithms labeled as our1 through our4. In the quantification accuracy experiment, we tested the accuracy with different numbers of groups $m$ and various sizes of data matrices. In the sensitivity experiment, we evaluated the impact of different numbers of groups $m$ on the results while maintaining the same compression ratio.
3.2 Experiment result
Accuracy Measurement
Figure 5: MAE on 8 subspaces.
Figure 6: MRE on 8 subspaces.
Figure 7: MSE on 8 subspaces.
Figure 8: MAE on 16 subspaces.
Figure 9: MRE on 16 subspaces.
Figure 10: MSE on 16 subspaces.
Figure 11: MAE on 32 subspaces.
Figure 12: MRE on 32 subspaces.
Figure 13: MSE on 32 subspaces.
Figure 14: MAE on 128 subspaces.
Figure 15: MRE on 128 subspaces.
Figure 16: MSE on 128 subspaces.
Figure 17: MAE on 1 iterations.
Figure 18: MRE on 1 iterations.
Figure 19: MSE on 1 iterations.
Figure 20: MAE on 2 iterations.
Figure 21: MRE on 2 iterations.
Figure 22: MSE on 2 iterations.
Figure 23: MAE on 3 iterations.
Figure 24: MRE on 3 iterations.
Figure 25: MSE on 3 iterations.
Figure 26: MAE on 4 iterations.
Figure 27: MRE on 4 iterations.
Figure 28: MSE on 4 iterations.
We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  7-  7, Figures  10-  10, Figures  13-  13, Figures  16-  16 respectively. For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
From the experimental results, we can derive four conclusions. Firstly, the higher the compression ratio, the less space is consumed, but the accuracy decreases. Secondly, the OPQ algorithm performs worse than others with the same compression rate because it needs to store the transformation matrix. Additionally, the PQ algorithm does not perform as well as our improved algorithm. Finally, the more iterations, the better the performance of our algorithm.
Parameter sensitivity analysis: We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance. The results for 1 to 4 iterations are shown in Figures  1919, Figures  $2$, and Figures  2525, respectively. Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).
4 Conclusion
In conclusion, our research addresses critical challenges in the optimization of large language models (LLMs) through the development of a novel algorithm, HERA. The growing complexity and scale of LLMs necessitate innovative approaches to enhance storage and computational efficiency, particularly in vector database management, weight quantization, and key-value (k-v) cache optimization.
HERA’s hierarchical approach to segmenting, compressing, and reorganizing the matrix dataset has proven to be effective in significantly reducing quantization error. By considering the distribution and magnitude of parameters, our method achieves superior performance compared to traditional uniform quantization techniques. The experimental results from our prototype system, implemented in Python, demonstrate that HERA can reduce the quantization error to 12.3% of the original, maintaining the same compression ratio.
Future work will explore further refinements to HERA, including its application to other types of application and broader datasets. Additionally, investigating the integration of HERA with other optimization techniques may yield even greater enhancements in model performance and efficiency. The promising results of this study encourage continued research and development in the quest for more effective and scalable solutions for LLM compression and optimization.

### v2

@@
 Abstract
-Large Language Models (LLMs) have significantly advanced natural language processing tasks such as machine translation, text generation, and sentiment analysis.
-However, their large size, often consisting of billions of parameters, poses challenges for storage, computation, and deployment, particularly in resource-constrained environments like mobile devices and edge computing platforms.
-Additionally, the key-value (k-v) cache used to speed up query processing requires substantial memory and storage, exacerbating these challenges.
-Vector databases have emerged as a crucial technology to efficiently manage and retrieve the high-dimensional vectors produced by LLMs, facilitating faster data access and reducing computational demands.
-Effective compression and quantization techniques are essential to address these challenges, as they reduce the memory footprint and computational requirements without significantly compromising performance.
-Traditional methods that uniformly map parameters to compressed spaces often fail to account for the uneven distribution of parameters, leading to considerable accuracy loss.
-Therefore, innovative approaches are needed to achieve better compression ratios while preserving model performance.
-In this work, we propose HERA, a novel algorithm that employs heuristic Element Replacement for compressing matrix.
-HERA systematically replaces elements within the model using heuristic methods, which simplifies the structure of the model and makes subsequent compression more effective.
-By hierarchically segmenting, compressing, and reorganizing the matrix dataset, our method can effectively reduce the quantization error to 12.3% of the original at the same compression ratio.
+The matrix quantization entails representing matrix elements in a more space-efficient form to reduce storage usage, with dequantization restoring the original matrix for use.
+We formulate the Quantization Error Minimization (QEM) problem as minimizing the distance between a matrix before and after quantization, under the condition that the quantized matrix occupies the same memory space.
+Matrix quantization is crucial in various applications, including Large Language Models (LLMs) weight quantization, vector databases, KV cache quantization, graph compression, and image compression.
+Recent advancements in LLMs, such as GPT-4 and BERT, have highlighted the importance of matrix compression due to the large size of parameters and KV cache, which are stored as matrices.
+We propose HETA (HETA) to address the QEM problem by leveraging the local orderliness of matrix elements, involving iterative element swapping to form a locally ordered matrix.
+This matrix is then grouped and quantized by columns.
+To enhance HETA, we introduce two optimizations: further quantizing residuals to reduce MSE, and using masking and batch processing to accelerate the algorithm.
+Experimental results demonstrate that HETA can effectively reduce MSE to 12.3% of the original value at the same compression ratio, outperforming the best baseline algorithms.
+Our contributions include the abstraction of the QEM problem, the design of the HETA algorithm, and the proposal of two optimizations to improve accuracy and speed.
 1 Introduction
+The matrix quantization entails representing the elements of a matrix in a more space-efficient form to reduce its storage usage.
+Dequantization, on the other hand, is the process during usage where the original matrix is restored from the quantized matrix using a restoration algorithm.
+We formulate the Quantization Error Minimization (QEM) problem as the task of minimizing the distance between a matrix before and after quantization in high-dimensional space, under the condition that the quantized matrix occupies the same space.
+Matrix quantization is widely employed across diverse applications, including Large Language Models (LLMs) weight quantization (Dettmers et al.
+(2024); Lin et al.
+(2024); Shao et al.
+(2023); Xiao et al.
+(2023)), vector database (Xu et al.
+(2018); Jegou et al.
+(2010)), LLM k-v cache quantization (Liu et al.
+(2024); Zhang et al.
+(2024); Hooper et al.
+(2024); Kawakibi Zuhri et al.
+(2024); Duanmu et al.
+(2024); Yue et al.
+(2024); Lee et al.
+(2024); Adnan et al.
+(2024)), graph compression (Brisaboa et al.
+(2009); Claude & Ladra (2011)), and image compression (Yu et al.
+(2018); Ning et al.
+(2016)).
+Specifically, the recent advancements in Large Language Models (LLMs) have made matrix compression even more critical.
 Large Language Models (LLMs) have revolutionized the field of natural language processing (NLP), enabling significant advancements in tasks such as machine translation, text generation, and sentiment analysis.
 These models, characterized by their large-scale neural network architectures and vast training datasets, have shown remarkable capabilities in understanding and generating human language.
-The advent of LLMs, such as OpenAI’s GPT-3 and BERT by Google, has pushed the boundaries of what machines can achieve in linguistic tasks, providing near-human performance in various applications  Brown et al.
-(2020); Devlin et al.
-(2018).
-The development of LLMs is rooted in the transformer architecture, which employs self-attention mechanisms to process and produce language with high contextual relevance.
-This architecture has replaced previous recurrent neural network (RNN) and long short-term memory (LSTM) models due to its efficiency and ability to handle long-range dependencies in text Vaswani et al.
-(2017).
-As a result, LLMs have become the cornerstone of modern NLP, driving innovations in areas such as automated customer service, content creation, and real-time language translation.
-Furthermore, the development and deployment of LLMs face significant challenges, particularly in optimizing vector database efficiency, LLM weight quantization, and key-value (k-v) cache optimization.
-First, as the size and complexity of LLMs increase, efficiently managing and retrieving vectors in large databases becomes critical.
-Optimizing vector database efficiency involves improving indexing and search algorithms to handle high-dimensional data, which is essential for tasks like semantic search and recommendation systems Johnson et al.
-(2019).
-Second, LLM weight quantization, which involves reducing the precision of model weights, is a crucial technique for making these models more storage and computation efficient without significantly degrading their performance.
-Quantization can drastically reduce the memory footprint and computational requirements, enabling the deployment of LLMs on resource-constrained devices.
-However, achieving optimal quantization while maintaining model accuracy remains a complex challenge that requires sophisticated algorithms and techniques Dettmers et al.
-(2021).
-Lastly, optimizing k-v cache efficiency is vital for speeding up inference times in LLMs.
-The k-v cache stores intermediate activations during the forward pass, which can be reused to avoid redundant computations.
-Efficient management and compression of these caches are essential to reduce latency and improve the throughput of LLMs, especially in real-time applications like chatbots and virtual assistants.
-Figure 1: The distribution of LLM parameter.
-To address the challenges of storage and computational efficiency in large language models (LLMs), it is crucial to develop techniques that optimize vector database efficiency, weight quantization, and key-value (k-v) cache optimization.
-Therefore, finding quantization algorithms that can efficiently compress and quickly decompress matrix data is essential.
-In practice, when compressing parameters, the distribution of parameters is often uneven.
-First, the distribution is not uniform, and second, the magnitudes of different data points vary significantly.
-For example, the distribution of Transformer neural network weights is shown in Figure  1.
-In Figure  1, different rows represent the 0th layer, 8th layer, 16th layer, and 24th layer.
-Different columns represent the attention layer’s q, k, v, and o layers, as well as the fully connected layer’s gate, up, and down layers.
-The distributions of different layers are not uniformly analyzed, and the distributions vary across different layers.
-Previous algorithms uniformly and linearly map parameters to another compressed space.
-Although this method is simple, it may not be effective and can result in significant accuracy loss because it does not take into account the original data distribution.
-It maps both densely populated and sparse regions to the quantized space in the same way.
-Besides, A lines of works have been proposed to optimize vector database efficiency Ge et al.
-(2013); Jegou et al.
-(2010); Kalantidis, Avrithis (2014), quantize the model weights Dettmers et al.
-(2022); Polino et al.
-(2018); Chmiel et al.
-(2020); Fan et al.
-(2020); Zafrir et al.
-(2019); Wu et al.
-(2022); LeCun et al.
-(1989) and optimizate kv cache Hooper et al.
-(2024); Cho et al.
-(2024).
-They also does not take into account the original data distribution.
-Therefore, a more reasonable approach is to quantize the model based on the distribution and magnitude of the parameters, stratifying and computing accordingly.
-In this work, we propose a novel algorithm named HERA to enhance the compression of large language models.
-The core idea of HERA is to group the relevant matrices and compress each group using clustering methods.
-Furthermore, HERA reorders the original dataset based on the distribution and magnitude of the parameters.
-This reordering is performed in multiple layers, with each layer featuring a new arrangement that optimizes the dataset’s distribution for compression.
-By leveraging these techniques, HERA effectively reduces data error and minimizes the discrepancy between estimated data and actual data.
-We implemented a prototype system using Python.
-Experimental results show that, by hierarchically segmenting, compressing, and reorganizing the matrix dataset, our method can effectively reduce the quantization error to 12.3% of the original at the same compression ratio.
-2 HERA
+The advent of LLMs, such as OpenAI’s GPT-4 (Achiam et al.
+(2023)), and BERT (Devlin et al.
+(2018)) by Google, has pushed the boundaries of what machines can achieve in linguistic tasks, providing near-human performance in various applications.
+The parameters and KV cache in Large Language Models are very large in size.
+For instance, GPT-2 contains 1.5 billion parameters (Solaiman et al.
+(2019)), whereas GPT-3 has expanded to 175 billion parameters (Brown et al.
+(2020)).
+Additionally, the KV cache accounts for over 30% of GPU memory during deployment, compared to the 65% occupied by the parameters (Kwon et al.
+(2023)).
+Both the parameters and KV cache are stored in the form of matrices.
+GPTQ (Frantar et al.
+(2022)) and SqueezeLLM (Kim et al.
+(2023)) directly address the QEM problem by treating it as the optimization objective for quantizing parameters and the KV cache.
+Therefore, the abstracted scientific problem of QEM is highly significant.
+There has been a growing literature on the matrix quantization.
+The first category of methods focuses on independently compressing the elements of a matrix for each specific scenario.
+Examples of such techniques include LLM.int8 (Ge et al.
+(2013)), Optimal Brain Damage (LeCun et al.
+(1989)), GPTQ (Frantar et al.
+(2022)), AWQ (Lin et al.
+(2023)), SmoothQuant (Xiao et al.
+(2023)), and OmniQuant (Shao et al.
+(2023)).
+The second category groups matrix elements by columns and then applies quantization to each group.
+Relevant works in this category include product quantization (PQ) (Jegou et al.
+(2010)), optimized product quantization (OPQ) (Ge et al.
+(2013)), locally optimized product quantization (LOPQ) (Kalantidis & Avrithis (2014)), and SqueezeLLM (Kim et al.
+(2023)).
+The first category of work can be summarized as utilizing outliers and the importance of elements in specific scenarios to improve the RTN (Round-To-Nearest)  [^1] RTN (Round-To-Nearest) is an algorithm that quantizes elements by rounding each value to its nearest representable level.
+(Gray & Neuhoff (1998) algorithm).
+The second category focuses on enhancements to PQ algorithms.
+The primary drawback of the first category is that it does not consider the correlation between elements, instead independently quantizing and storing elements.
+Conversely, the second category fails to account for the relative order of elements.
+In our research, we observed that the order of elements has a significant impact on the quantization outcome.
+Intuitively, consider a matrix $\begin{bmatrix}1&2\\ 2&1\end{bmatrix}$.
+To losslessly quantize this matrix (with MSE=0), two vectors, $[1,2]$ and $[2,1]$, are required.
+However, if the matrix is reordered to $\begin{bmatrix}1&2\\ 1&2\end{bmatrix}$, it can be losslessly quantized using just a single vector, $[1,2]$.
+Using these ideas, we propose HETA (HETA) to address the QEM problem in matrix quantization.
+The core idea of HETA is to leverage the local orderliness of matrix elements to optimize the QEM problem.
+The design involves swapping adjacent elements of the matrix to form a new, locally ordered matrix.
+To cover a broader range, we can perform further element swapping on the initially locally ordered matrix, similar to the approach of receptive field used in convolutional neural networks (LeCun et al.
+(1998)).
+This step can be repeated for multiple iterations.
+The newly ordered matrix is then grouped by columns, with each group being quantized.
+Additionally, we propose two optimization algorithms based on the basic HETA.
+First, the residuals between the original and quantized matrices can be further quantized to enhance the accuracy of the results.
+Second, since the iterative swapping operations can slow down the algorithm, we introduce optimization techniques based on masking and batch processing to accelerate the HETA.
+We evaluate HETA on multiple datasets.
+Experimental results demonstrate that our method can effectively reduce the MSE to 12.3% of the original value at the same compression ratio, compared to the best baseline algorithm, which achieves an MSE of XXXX%.
+We summarize our contributions below.
+  • Abstracted a problem: We abstracted the Quantization Error Minimization (QEM) problems from various application scenarios.
+  • Designed an algorithm: We developed the HETA (HETA) algorithm, leveraging the concept of local orderliness to optimize the QEM problem.
+  • Proposed two optimizations: We introduced the use of residuals to reduce MSE and employed batch processing techniques to accelerate HETA algorithm.
+2 PROBLEM SETTING
+In this section, we formally define the Quantization Error Minimization (QEM) problem and analyze the impact and feasibility of rearranging matrix elements prior to quantization.
+QEM Problem: we denote the matrix to be quantized as $X$, the quantized matrix as $X^{q}$, and the matrix obtained by dequantizing $X^{q}$ as $X^{\prime}$.
+Both $X$ and $X^{\prime}$ are $n\times d$ dimensional, but the dimensions of $X^{q}$ depend on the quantization algorithm.
+The elements of these matrices are denoted as $x_{(i,j)}$, $x^{q}_{(i,j)}$, and $x^{\prime}_{(i,j)}$, respectively.
+The memory occupied by a matrix is denoted as memory(), and the memory constraint as mem_constrain.
+Definition 1 (Quantization Error Minimization (QEM) Problem).
+The Quantization Error Minimization (QEM) problem is defined as the task of minimizing the combined objective of the Mean Squared Error (MSE) and the memory size of the quantized matrix:
+$\text{minimize}\;\text{MSE}(X,X^{\prime})+\lambda\cdot\text{memory}(X^{q}),$
+where
+$\text{MSE}(X,X^{\prime})=\frac{1}{n\cdot d}\sum_{i,j}\left(x_{(i,j)}-x^{\prime}_{(i,j)}\right)^{2},$
+$\lambda$ is a regularization parameter that balances the trade-off between minimizing the MSE and the memory usage of the quantized matrix.
+Subject to the condition that the quantized matrix occupies the same space:
+[…]
+Rearranging matrix elements: Our idea is to rearrange the matrix elements before quantization.
+Let the quantization algorithm be denoted as quant() and the rearrangement algorithm as rearrange().
+Previously, the quantized matrix was obtained as $X^{q}=\text{quant}(X)$, whereas in our rearrangement approach, it is obtained as $X^{q}=\text{quant}(\text{rearrange}(X))$.
+Therefore, the rearrangement algorithm only affects the quantization if the quantization method is order-sensitive.For instance, order-insensitive algorithms like RTN do not benefit from rearrangement because RTN processes each element independently, while order-sensitive algorithms like PQ can be optimized through rearrangement.
+However, finding the optimal arrangement requires exploring a large search space.
+Specifically, for an $n\times d$ matrix, there are $(n!)^{d}$ possible rearrangements.
+Therefore, searching the entire space is impractical, and we need heuristic algorithms to prioritize exploring more promising regions of the search space.
+3 METHODS
+The core idea of HETA (HETA) is to leverage the local orderliness of matrix elements by rearranging adjacent elements to optimize the matrix quantization algorithm.
+Additionally, multiple iterations are used to expand the algorithm’s swapping field.
+We begin by describing the basic version of this idea and then propose two optimizations for the basic algorithm.
 3.1 Basic algorithm
-We borrow the idea from product quantization algorithm and k-means algorithm to compress the model parameter.
-The algorithm is divided into four steps, the first three steps is quantization process as shown in Figure  2, the last two steps is the dequantization process as show in Figure  3.
-The detailed process is described below.
-Algorithm 1 Product Quantization Algorithm
-1:$X\in\mathbb{R}^{N\times D}$triangleright Input matrix
-2:Quantization Process:
-3:Step 1: Space Partitioning
-4:for each vector $x_{i}$ in $X$ do
-5:     Divide $x_{i}$ into $M$ low-dimensional subspaces $x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{M}$
-6:end for
-7:Step 2: Codebook Calculation
-8:for each subspace $j$ from 1 to $M$ do
-9:     Apply k-means algorithm to subspace $j$ to get $k$ cluster centers
-10:     Create codebook $C_{j}$ for subspace $j$ with cluster centers
-11:end for
-12:Step 3: Quantization
-13:for each vector $x_{i}$ in $X$ do
-14:     for each subspace $j$ from 1 to $M$ do
-15:         Map $x_{i}^{j}$ to the nearest cluster center in $C_{j}$
-16:     end for
-17:end for
-18:Dequantization Process:
-19:Step 4: Codebook Reconstruction
-20:Construct a global codebook $C$ that contains all possible combinations of cluster centers from $C_{1},C_{2},\ldots,C_{M}$
-21:Step 5: Dequantization
-22:for each quantized vector $q_{i}$ in the quantized dataset do
-23:     Restore $q_{i}$ to its original high-dimensional space using the global codebook $C$
-24:end for
-  1. 1.
-    Space Partitioning: Divide the high-dimensional space into a series of low-dimensional subspaces.
-  2. 2.
-    Codebook calculation: Use k-means algorithm to calculate codebook of low-dimensional subspaces.
-  3. 3.
-    Quantization: Perform independent quantization within each low-dimensional subspace, mapping data points to the nearest cluster center.
-  4. 4.
-    Codebook Reconstruction: Reconstruct the data using a codebook that contains all possible cluster centers.
-  5. 5.
-    Dequantization: Restore the vectors in the original high-dimensional space based on the vectors in the quantized low-dimensional subspaces.
-Figure 2: The Quantization Process of HERA.
-Figure 3: The Dequantization Process of HERA.
-In data matrices where the distribution of data is uneven across different columns, directly applying the k-means algorithm for clustering may yield suboptimal results.
-This unevenness can distort the clustering process, as k-means assumes that all features contribute equally to the distance calculations.
-The differing distributions across columns mean that some features may disproportionately influence the clustering outcome, leading to biased clusters that do not accurately reflect the true structure of the data.
-To mitigate these effects, it is crucial to preprocess the data appropriately.
-One common approach is to normalize the data so that each column contributes equally to the distance metric used by k-means.
-Standard normalization techniques, such as z-score normalization or min-max scaling, can be employed to transform the data into a more uniform distribution.
-By ensuring that all features are on a comparable scale, the k-means algorithm can perform more effectively, producing clusters that better represent the underlying patterns in the data.
-Algorithm 2 Product Quantization with HERA heuristics
-1:Input matrix: $X\in\mathbb{R}^{N\times D}$, number of iteration in HERA: $iter$.
-2:Quantized feature matrix and feature map
-3:for each element $x_{i}\in X$ do
-4:     Rearrange $X$ into pairs: $X_{\text{pairs}}\in\mathbb{R}^{N\times D/2}$
-5:     for each dimension $j$ do
-6:         $x_{pair_{i,j}}=(x_{2i,j},x_{2i+1,j})\in X_{\text{pairs}}$
+The steps and design of the HETA algorithm are illustrated in Figure 1. The figure provides a conceptual overview of the HETA (HETA) algorithm.
+On the left side of the figure, the process of rearranging matrix elements through multiple iterations in the HETA is depicted.
+This iterative procedure involves comparing adjacent elements and ensuring that adjacent elements are not placed in the same matrix.
+This separation of adjacent elements into different matrices is conceptually similar to quantum entanglement, hence the name ”HETA.” The HETA algorithm performs multiple iterations, with each iteration comparing adjacent elements and splitting them based on their values.
+This approach is based on two observations.
+First, the local orderliness of the matrix can enhance the regularity of matrix elements, thereby improving compression efficiency.
+Second, multiple iterations allow the orderliness to cover a larger number of elements.
+For instance, adjacent elements after the first iteration might be non-adjacent in the original matrix.
+This iterative process, similar to that of convolutional neural networks, extends the ”receptive field” to cover more elements.
+On the right side of the figure, the HETA compression process is shown.
+After the elements have been rearranged, the algorithm groups the sub-matrices, followed by clustering to determine the centroids.
+These centroids are used to quantize the matrix.
+The final quantized matrix, codebook, and indicator maps are the outputs of the HETA algorithm.
+This divide-and-conquer compression method is based on an observation: by splitting the matrix, the expressive power can be increased under a fixed storage constraint.
+For example, if the matrix is divided into $m$ blocks, each with $c$ centroids, the entire matrix can express $c^{m}$ centroids while only storing $c\times m$ centroids.
+Without splitting, storing $c^{m}$ centroids would be required.
+3.2 HETA Quantization
+The proposed HETA Algorithm is designed to optimize the quantization of a matrix $X\in\mathbb{R}^{n\times d}$ by leveraging local orderliness.
+The algorithm proceeds in four key steps, as shown in Algorithm 2:
+Step 1: Initial Partitioning.
+The algorithm starts by comparing each adjacent pair of elements $(x_{i,j},x_{i,j+1})$ in the input matrix $X$.
+Based on the comparison of these elements, the smaller value is placed into matrix $S$ and the larger into matrix $L$.
+Concurrently, an Indicator Map $I$ is generated, where a 0 denotes that the smaller element is on the left and a 1 denotes it is on the right.
+This process results in the creation of two matrices, $S$ and $L$, which collectively represent the initial partitioning of the original matrix $X$.
+Step 2: Recursive Partitioning.
+Following the initial partitioning, the algorithm undertakes a recursive partitioning process.
+Starting from $k=1$, the matrices $S^{i}_{k}$ and $L^{i}_{k}$ [^2] Since each level has multiple $S$ and $L$ matrices, we use the subscript to denote the iteration number and the superscript to denote the sequence within each iteration.
+derived from the previous iteration (for $i=1,2,\dots,2^{k-1}$) are further partitioned based on their sizes.
+In each iteration, a new Indicator Map $I_{k+1}$ is generated and then merged with the existing Indicator Maps.
+This recursive process continues until the predefined number of iterations $l$ is reached.
+The resulting Indicator Maps $I_{1},I_{2},\dots,I_{l}$ are stored in the set $IM$.
+The final matrices $S_{l}$ and $L_{l}$ are combined to form a new matrix $X^{*}$, which represents a locally ordered version of the original matrix.
+Step 3: Subspace Grouping.
+After constructing the matrix $X^{*}$, it is divided into subspaces $\{G_{i}\}$, where $i=1,2,\dots,m$.
+This division is based on the local ordering of elements, which facilitates more effective clustering and quantization in the subsequent steps.
+Step 4: Clustering and Quantization.
+In the final step, each subspace $G_{k}$ (for $k=1,2,\dots,m$) is clustered to determine a set of centroids $C_{k}=\{\mathbf{c}_{j}\}$.
+Each vector $\mathbf{g_{i}}$ within the subspace is then quantized by assigning it to the nearest centroid.
+This process results in the quantized matrix $X^{q}$.
+The output of the algorithm includes the quantized matrix $X^{q}$, the codebook $\mathbf{C}$ containing the centroids, and the Indicator Maps $IM$.
+Algorithm 1 HETA Quantization Algorithm
+1:Input: Matrix $X\in\mathbb{R}^{n\times d}$, $m$ is the number of subspaces for quantization, $l$ is the number of HETA iterations.
+2:Output: Quantized matrix $X^{q}=\{X^{q}_{i,k}\}$ where $i=1,\dots,d$ and $k=1,\dots,m$, Codebook $\mathbf{C}=\{C_{i}\}$ where $i=1,\dots,m$, and Indicator Maps $IM=\{I_{i}\}$ where $i=1,\dots,l$
+3:Initialize $IM$ and $X^{q}$ as empty
+4:Step 1: Initial Partitioning
+5:for each adjacent pair $(x_{i,j},x_{i,j+1})$ in $X$ do
+6:     if $x_{i,j}>x_{i,j+1}$ then
+7:         $S_{i,j}\leftarrow x_{i,j+1}$, $L_{i,j}\leftarrow x_{i,j}$
+8:         $I\left(\frac{i}{2},j\right)\leftarrow 0$
+9:     else
+10:         $S_{i,j}\leftarrow x_{i,j}$, $L_{i,j}\leftarrow x_{i,j+1}$
+11:         $I\left(\frac{i}{2},j\right)\leftarrow 1$
+12:     end if
+13:end for
+14:Step 2: Recursive Partitioning
+15:$k\leftarrow 1$
+16:while $k\neq l$ do
+17:     for $i=1$ to $2^{k-1}$ do
+18:         Apply size-based partitioning to $S^{i}_{k}$ and $L^{i}_{k}$
+19:         Generate and merge new Indicator Maps into $I_{k+1}$ for each partition
+20:     end for
+21:     $k\leftarrow k+1$
+22:end while
+23:Store $I_{1},I_{2},\dots,I_{l}$ into $IM$
+24:$X^{*}\leftarrow S^{1}_{l}\cup L^{1}_{l}\cup S^{2}_{l}\cup L^{2}_{l}\cup\dots\cup S^{2^{l-1}}_{l}\cup L^{2^{l-1}}_{l}$triangleright Combine the locally ordered parts from the final layer $l$
+25:Step 3: Subspace Grouping
+26:Group the matrix $X^{*}$ into subspaces $\{G_{i}\}$ where $i=1,2,\dots,m$
+27:Step 4: Clustering and Quantization
+28:for each vector $\mathbf{g_{i}}$ in $G_{k}$ (for $k=1,2,\dots,m$do
+29:     Apply clustering to find centroids $C_{k}=\{\mathbf{c}_{j}\}$
+30:     $X^{q}_{i,k}\leftarrow\arg\min_{\mathbf{c}_{j}}\|\mathbf{g_{i}}-\mathbf{c}_{j}\|$triangleright Quantize to nearest centroid within group
+31:end for
+32:Return Quantized matrix $X^{q}$, Codebook $\mathbf{C}$, and Indicator Maps $IM$
+3.2.1 HETA Dequantization
+The dequantization process, as described by the HETA Dequantization Algorithm, aims to reconstruct the original matrix $X^{\prime}$ from its quantized version $X^{q}$.
+This process involves several key steps, as shown in Algorithm 2:
+Step 1: Dequantization Using Codebook.
+The first step involves using the codebook $\mathbf{C}$ to directly retrieve the original vectors corresponding to each quantized vector $\mathbf{g_{i}^{q}}$ in the matrix $X^{q}$.
+For each vector $\mathbf{g_{i}^{q}}$ in $X^{q}_{i,k}$, the algorithm assigns $\mathbf{g_{i}^{\prime}}$ to the centroid $\mathbf{c}_{j}$ from the codebook, where $j$ is the index corresponding to the quantized vector.
+This step effectively reverses the quantization by mapping the quantized vectors back to their original forms using the codebook entries.
+Step 2: Reverse Subspace Grouping.
+After mapping, the algorithm combines the dequantized subspaces $\{G_{i}^{\prime}\}$, where $i=1,2,\dots,m$, into a single matrix $X^{\prime*}$.
+Each subspace $G_{i}^{\prime}$ contains the vectors $\mathbf{g_{i}^{\prime}}$ corresponding to the original matrix entries.
+This step reconstructs the subspaces into an uniform matrix that is a preliminary version of the original matrix $X^{\prime}$.
+Step 3: Reverse Recursive Partitioning and Final Reconstruction.
+The final step involves reversing the recursive partitioning process that was applied during quantization.
+Starting from the last iteration $k=l$, the algorithm merges the matrices $S_{k}$ and $L_{k}$ based on the Indicator Maps $I_{k}$ and the matrix $X^{\prime*}$.
+This merging process is iteratively applied from $k=l$ down to $k=1$, updating $X^{\prime*}$ at each step to gradually reconstruct the original matrix structure.
+The final output of the algorithm is the reconstructed matrix $X^{\prime}\in\mathbb{R}^{n\times d}$, which approximates the original matrix before quantization.
+Algorithm 2 HETA Dequantization Algorithm
+1:Input: Quantized matrix $X^{q}=\{X^{q}_{i,k}\}$, where $i=1,\dots,d$ and $k=1,\dots,m$; Codebook $\mathbf{C}=\{C_{i}\}$, where $i=1,\dots,m$; and Indicator Maps $IM=\{I_{i}\}$, where $i=1,\dots,l$
+2:Output: Reconstructed matrix $X^{\prime}\in\mathbb{R}^{n\times d}$
+3:Initialize $X^{\prime}$ as empty
+4:Step 1: Dequantization Using Codebook
+5:for each vector $\mathbf{g_{i}^{q}}$ in $X^{q}_{i,k}$ do
+6:     $\mathbf{g_{i}^{\prime}}\leftarrow\mathbf{c}_{j}$, where $j$ corresponds to the codebook entry for $\mathbf{g_{i}^{q}}$triangleright Direct lookup from codebook
 7:end for
-8:end for
-9:for each pair $x_{pair_{i,j}}$ in $X_{\text{pairs}}$ do
-10:     if $x_{pair_{i,j}}[1]<x_{pair_{i,j}}[2]$ then
-11:         $X[iter]_{small_{i,j}}\leftarrow x_{pair_{i,j}}[1]$
-12:         $X[iter]_{big_{i,j}}\leftarrow x_{pair_{i,j}}[2]$
-13:         $fm[iter]_{i,j}\leftarrow 1$
-14:     else
-15:         $X[iter]_{small_{i,j}}\leftarrow x_{pair_{i,j}}[2]$
-16:         $X[iter]_{big_{i,j}}\leftarrow x_{pair_{i,j}}[1]$
-17:         $fm[iter]_{i,j}\leftarrow 0$
-18:     end if
-19:end for
-20:for $X_{small}$ and $X_{large}$ do
-21:     code[$iter$], codebook [$iter$] $\leftarrow$ PQ($X[iter]_{small}$)
-22:     code[$iter$], codebook[$iter$] $\leftarrow$ PQ($X[iter]_{large}$)
-23:end for
-24:$iter$leftarrow $iter$ 1
-25:if $iter$neqthen
-26:     Goto line 1
-27:end if
-However, these transformations require specific parameters for each column and involve complex calculations, which can be computationally intensive and demand significant storage resources.
-To address this, we employ a reordering method that permutations the elements of the matrix based on their relative sizes.
-This approach achieves a more uniform distribution, enabling more effective application of compression operations.
-Algorithm 3 Dequantization of Product Quantization with HERA heuristics
-1:Quantized feature matrices: $code_{small}$, $code_{big}$, codebooks: $codebook_{small}$, $codebook_{big}$, feature maps: $fm$, number of iterations in HERA: $iter$
-2:Reconstructed matrix: $\hat{X}\in\mathbb{R}^{N\times D}$
-3:Initialize $\hat{X}_{pairs}\in\mathbb{R}^{N\times D/2}$
-4:for each iteration from $1$ to $iter$ do
-5:     for each pair $x_{pair_{i,j}}$ in $\hat{X}_{pairs}$ do
-6:         $\hat{X}_{small_{i,j}}\leftarrow\text{decode}(code_{small}[iter],codebook_{small}[iter])$
-7:         $\hat{X}_{big_{i,j}}\leftarrow\text{decode}(code_{big}[iter],codebook_{big}[iter])$
-8:     end for
-9:     for each pair $x_{pair_{i,j}}$ in $\hat{X}_{pairs}$ do
-10:         if $fm[iter]_{i,j}=1$ then
-11:              $x_{pair_{i,j}}[1]\leftarrow\hat{X}_{small_{i,j}}$
-12:              $x_{pair_{i,j}}[2]\leftarrow\hat{X}_{big_{i,j}}$
-13:         else
-14:              $x_{pair_{i,j}}[2]\leftarrow\hat{X}_{small_{i,j}}$
-15:              $x_{pair_{i,j}}[1]\leftarrow\hat{X}_{big_{i,j}}$
-16:         end if
-17:     end for
-18:end for
-19:for each dimension $j$ do
-20:     for each element $i$ in $\hat{X}_{pairs}$ do
-21:         $\hat{x}_{2i,j}\leftarrow x_{pair_{i,j}}[1]$
-22:         $\hat{x}_{2i+1,j}\leftarrow x_{pair_{i,j}}[2]$
-23:     end for
-24:end for
-The purpose of permutation is to explore all possible reordering methods within the search space. For an $N\times D$ matrix, there are $(N!)^{D}$ possible permutations, creating a vast search space.
-To manage this, we employ heuristic algorithms to reduce the search space and optimize the algorithm’s complexity.
-Heuristic algorithms provide approximate solutions by focusing on the most promising areas of the search space, rather than exhaustively evaluating all possible permutations.
-These methods significantly decrease computational requirements while still delivering high-quality solutions.
-By adopting this heuristic reordering strategy, we aim to achieve a more balanced distribution of data across the matrix.
-This not only simplifies the preprocessing steps but also enhances the efficiency of subsequent clustering and compression processes.
-Ultimately, this method helps overcome the challenges posed by uneven data distributions, leading to more accurate and resource-efficient clustering outcomes.
-2.2 HERA heuristics
-To improve quantization and avoid uneven distribution, HERA first perceives the distribution of data sizes, then rearranges the elements according to the distribution, and finally quantizes the rearranged matrix.
-2.2.1 HERA Design
-We observe that compressing matrix elements of similar sizes as a single column (group) yields better results.HERA uses a low-cost method to locally adjust element arrangement based on data sizes.
-We first introude the HERA design.
-The design of HERA heuristics is shown in Figure  4.
-Figure 4: HERA heuristics.
-HERA groups every two elements in the matrix together, rearranges the matrix elements based on the size of adjacent elements, and then quantizes the rearranged results.
-Quantization: The quantization process is illustrated in the algorithm 2, the quantization process consists of multiple steps, which involve rearranging the input data, comparing elements, and quantizing both smaller and larger value sets.
-The algorithm begins by rearranging the input matrix $X\in\mathbb{R}^{N\times D}$ into pairs (lines 1 in Algorithm 2).
-Each element $x_{i}\in X$ is grouped into $X_{\text{pairs}}\in\mathbb{R}^{N\times D/2}$ by combining every two consecutive elements along each dimension.
-For each dimension $j$, the pairs are formed as $(x_{2i,j},x_{2i+1,j})$ (lines 2–4 in Algorithm 2).
-Next, each pair $x_{pair_{i,j}}$ is compared to separate the smaller and larger values.
-If the first element of the pair is smaller, it is assigned to $X_{small}$ and the second to $X_{big}$, with a feature map $fm$ recording the comparison result as 1 (lines 5–8 in Algorithm 2).
-Otherwise, the assignment is reversed, and the feature map records a 0 (lines 9–12 in Algorithm 2).
-The quantization process proceeds by applying the PQ algorithm to both $X_{small}$ and $X_{big}$.
-This process is iteratively repeated for a predefined number of iterations ($iter$), refining the quantization results with each iteration (lines 13–16 in Algorithm 2).
-Dequantization:The Decompression process consists of multiple steps, namely decoding the quantized values, reconstructing the pairs, and restoring the original matrix dimensions.
-The algorithm begins by initializing the reconstructed matrix $\hat{X}_{pairs}\in\mathbb{R}^{N\times D/2}$ (line 1 in Algorithm 3).
-For each iteration, the algorithm decodes the quantized values from the codebooks for both small and large pairs (lines 2–4 in Algorithm 3).
-The decoded values are stored in $\hat{X}_{small}$ and $\hat{X}_{big}$ respectively.
-Next, for each pair in $\hat{X}_{pairs}$, the algorithm reconstructs the pairs based on the feature map $fm[iter]$.
-If the feature map value is 1, the first element of the pair is assigned the decoded small value and the second element the decoded big value (lines 5–8 in Algorithm 3).
-If the feature map value is 0, the assignment is reversed (lines 9–12 in Algorithm 3).
-Finally, the pairs are expanded back into the original dimensions to form the reconstructed matrix $\hat{X}$ (lines 13–17 in Algorithm 3).
+8:Step 2: Reverse Subspace Grouping
+9:Combine subspaces $\{G_{i}^{\prime}\}$, where $i=1,2,\dots,m$, $G_{i}^{\prime}=\{\mathbf{g_{i}^{\prime}}\}$ for $i=1$ to $n$, to form $X^{\prime*}$
+10:Step 3: Reverse Recursive Partitioning
+11:$k\leftarrow l$
+12:while $k\neq 0$ do
+13:     for $i=1$ to $2^{k-1}$ do
+14:         Merge $S^{i}_{k}$ and $L^{i}_{k}$ based on $I_{k}$ and $X^{\prime*}$, then update the results into $X^{\prime*}$
+15:     end for
+16:     $k\leftarrow k-1$
+17:end while
+18:Return Reconstructed matrix $X^{\prime}$
+3.3 Optimizations
+In this section, we propose three optimizations: Residual Optimization (RO), Batch Processing Optimization (BPO), and Codebook Quantization Optimization (CQO).
+RO takes advantage of the fact that, after applying HETA, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE.
+On the other hand, BPO utilizes batch processing techniques to enhance processing speed.
+Since the codebook used in the HETA algorithm occupies significant space, CQO further quantizes the codebook to improve space efficiency.
+Table 1: Range Ratios, and Codebook(CB) Proportions in different compression ratios.
+Compression Ratio 4.0 8.0 12.0 16.0
+Range Ratio (%) 1.6 4.0 9.2 29.7
+CB Proportion (%) 87.5 78.1 71.8 62.5
+3.3.1 Residual Optimization
+Residual Optimization (RO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the HETA algorithm.
+As shown in Table 1, our experiments indicate that at a compression ratio of 16, the data range of the quantized matrix is reduced to 29.7% of the original matrix’s range.
+Moreover, with a compression ratio of 4, the data range of the quantized matrix is reduced to approximately 1.6% of the original matrix’s range.
+Matrix quantization algorithms are more efficient at compressing matrix with a smaller data range; therefore, we propose Residual Optimization (RO).
+The RO algorithm, as depicted in Figure 1, begins with the quantization of the original matrix to produce a quantized matrix.
+This matrix is then dequantized to yield a dequantized matrix, which is subtracted from the original matrix to form a residual matrix.
+The residual matrix typically exhibits a significantly reduced data range, making it easier to compress.
+This residual matrix is then quantized using a chosen matrix quantization algorithm, such as RTN or HETA.
+During the dequantization process, the quantized residual is dequantized, and this dequantized residual is added back to the initially dequantized matrix to produce a revised dequantized matrix.
+Figure 1: Residual Optimization Algorithm.
+3.3.2 Batch Processing Optimization
+The key insight behind (Batch Processing Optimization) BPO is that by organizing and processing data in larger chunks or batches and enhancing the efficiency of the HETA algorithm through batch processing and vectorization operations, the algorithm can more efficiently determine the relationships between elements, such as identifying smaller and larger parts or setting indicator maps.
+In the original process, elements of the matrix are compared and sorted individually, leading to a sequential, element-by-element approach that can become computationally intensive as the matrix size increases.
+This method is not optimal for large-scale matrices.
+BPO introduces a more advanced technique by processing matrix elements in parallel batches rather than sequentially.
+This parallelization allows the algorithm to handle multiple comparisons and sorting operations simultaneously, rather than one at a time.
+By doing so, it reduces the overall computational complexity and accelerates the entire quantization process.
+3.3.3 Codebook Quantization Optimization (CQO)
+The codebook occupies a significant portion of the space in the HETA algorithm.
+As shown in Table 1, our experiments indicate that when the compression ratio is 16, the codebook occupies 62.5% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space.
+Therefore, compressing the codebook can significantly reduce the space used by the algorithm.
+Our optimization mainly employs the RTN method to compress the codebook.
+4 Theory
 5 Experiments
 5.1 Experiment setup
-Dataset: The datasets used in our experiments were generated with matrices of size $N\times D$, where $N$ is the number of samples and $D$ is the dimensionality of each sample.
-We employed one types of normal distribution to create these datasets.
-Normal Distribution Dataset: For the normal distribution dataset, each element of the matrix was drawn from a truncated normal distribution with a mean of 0.5 and a standard deviation of 0.16.
-The distribution was truncated to the open interval $(0,1)$ to ensure all values remain within this range.
-$\mathbf{X}_{ij}\sim\mathcal{TN}(0.5,0.16,0,1),\quad\forall i\in\{1,\ldots,N\},\quad\forall j\in\{1,\ldots,D\}$
-Here, $\mathcal{TN}(\mu,\sigma,a,b)$ denotes the truncated normal distribution with mean $\mu$, standard deviation $\sigma$, truncated to the interval $(a,b)$.
+Dataset: To evaluate the effectiveness of our proposed algorithm, we conducted experiments using one synthetic dataset and four real-world datasets: a synthetic normal distribution dataset, LLM weight dataset, KV cache dataset, Gist dataset, and Wiki dataset.
+Due to space constraints, we present the results for the synthetic normal distribution dataset and the LLM weight dataset in the main text.
+The results for the other datasets are provided in the appendix (Section A.1).
+Below, we provide a detailed description of each dataset.
+(1) Synthetic normal distribution dataset: This dataset was generated by drawing each element of the matrix from a truncated normal distribution with a mean of 0.5 and a standard deviation of 0.16.
+Additionally, one out of every ten thousand elements was replaced with an outlier value, randomly chosen between -100 and 100.
+(2) LLM weight dataset: This dataset comprises weight matrices extracted from the large language model (LLM) LLaMA2 (Touvron et al. (2023)).
+(3) KV cache dataset: The KV cache dataset is derived from the key-value pairs stored in the cache during inference in KV Quant (Hooper et al.
+(2024)).
+The key and value are stored in matrices, respectively.
+(4) Gist dataset: The Gist dataset consists of feature vectors used for image recognition and retrieval tasks.
+This dataset is particularly useful for evaluating the algorithm’s effectiveness in handling high-dimensional and dense data representations.
+(Jegou et al.
+(2010)) (Oliva & Torralba (2001))
+(5) Wiki dataset: The Wiki dataset contains document embeddings derived from Wikipedia articles. This dataset tests the algorithm’s ability to compress and process large-scale text data efficiently.
 Platform and implementation: We conducted our algorithm evaluations on a high-performance server equipped with an Intel Core i9-10980XE processor, featuring 18 cores and 36 threads, operating at a base frequency of 3.00 GHz.
@@
 In each repetition, matrices of the same size and from the same distribution were generated using different seeds.
+Metrics:
+Compared algorithms:
 Comparison of Algorithms and Parameter Selection
@@
 In the sensitivity experiment, we evaluated the impact of different numbers of groups $m$ on the results while maintaining the same compression ratio.
-3.2 Experiment result
 Accuracy Measurement
@@
 Figure 25: MSE on 4 iterations.
-We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  7-  7, Figures  10-  10, Figures  13-  13, Figures  16-  16 respectively.
+We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  4-  4, Figures  7-  7, Figures  10-  10, Figures  13-  13 respectively.
 For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
@@
 Finally, the more iterations, the better the performance of our algorithm.
-Parameter sensitivity analysis: We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance.
-The results for 1 to 4 iterations are shown in Figures  1919, Figures  $2$, and Figures  2525, respectively.
+We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance.
+The results for 1 to 4 iterations are shown in Figures  1616, Figures  1919, and Figures  $2$, respectively.
 Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).
@@
 The promising results of this study encourage continued research and development in the quest for more effective and scalable solutions for LLM compression and optimization.
+Appendix A Appendix
+A.1 Experimental Results on Additional Datasets

### v3

@@
-HERA: High-efficiency Matrix Compression via Element Replacement
+Quantum Entanglement Trees: Optimizing Quantized Matrix Quantization via Element Replacement and Residual Clustering
 Abstract
@@
 Recent advancements in LLMs, such as GPT-4 and BERT, have highlighted the importance of matrix compression due to the large size of parameters and KV cache, which are stored as matrices.
-We propose HETA (HETA) to address the QEM problem by leveraging the local orderliness of matrix elements, involving iterative element swapping to form a locally ordered matrix.
+We propose Quantum Entanglement Trees (QET) to address the QEM problem by leveraging the local orderliness of matrix elements, involving iterative element swapping to form a locally ordered matrix.
 This matrix is then grouped and quantized by columns.
-To enhance HETA, we introduce two optimizations: further quantizing residuals to reduce MSE, and using masking and batch processing to accelerate the algorithm.
-Experimental results demonstrate that HETA can effectively reduce MSE to 12.3% of the original value at the same compression ratio, outperforming the best baseline algorithms.
-Our contributions include the abstraction of the QEM problem, the design of the HETA algorithm, and the proposal of two optimizations to improve accuracy and speed.
+To enhance QET, we introduce two optimizations: further quantizing residuals to reduce MSE, and using masking and batch processing to accelerate the algorithm.
+Experimental results demonstrate that QET can effectively reduce MSE to 12.3% of the original value at the same compression ratio, outperforming the best baseline algorithms.
+Our contributions include the abstraction of the QEM problem, the design of the QET algorithm, and the proposal of two optimizations to improve accuracy and speed.
 1 Introduction
@@
 The first category of work can be summarized as utilizing outliers and the importance of elements in specific scenarios to improve the RTN (Round-To-Nearest)  [^1] RTN (Round-To-Nearest) is an algorithm that quantizes elements by rounding each value to its nearest representable level.
-(Gray & Neuhoff (1998) algorithm).
+(Gray & Neuhoff (1998)) algorithm.
 The second category focuses on enhancements to PQ algorithms.
@@
 However, if the matrix is reordered to $\begin{bmatrix}1&2\\ 1&2\end{bmatrix}$, it can be losslessly quantized using just a single vector, $[1,2]$.
-Using these ideas, we propose HETA (HETA) to address the QEM problem in matrix quantization.
-The core idea of HETA is to leverage the local orderliness of matrix elements to optimize the QEM problem.
+Using these ideas, we propose Quantum Entanglement Trees (QET) to address the QEM problem in matrix quantization.
+The core idea of QET is to leverage the local orderliness of matrix elements to optimize the QEM problem.
 The design involves swapping adjacent elements of the matrix to form a new, locally ordered matrix.
@@
 The newly ordered matrix is then grouped by columns, with each group being quantized.
-Additionally, we propose two optimization algorithms based on the basic HETA.
+Additionally, we propose two optimization algorithms based on the basic QET.
 First, the residuals between the original and quantized matrices can be further quantized to enhance the accuracy of the results.
-Second, since the iterative swapping operations can slow down the algorithm, we introduce optimization techniques based on masking and batch processing to accelerate the HETA.
-We evaluate HETA on multiple datasets.
+Second, since the iterative swapping operations can slow down the algorithm, we introduce optimization techniques based on masking and batch processing to accelerate the QET.
+We evaluate QET on multiple datasets.
 Experimental results demonstrate that our method can effectively reduce the MSE to 12.3% of the original value at the same compression ratio, compared to the best baseline algorithm, which achieves an MSE of XXXX%.
@@
   • Abstracted a problem: We abstracted the Quantization Error Minimization (QEM) problems from various application scenarios.
-  • Designed an algorithm: We developed the HETA (HETA) algorithm, leveraging the concept of local orderliness to optimize the QEM problem.
-  • Proposed two optimizations: We introduced the use of residuals to reduce MSE and employed batch processing techniques to accelerate HETA algorithm.
+  • Designed an algorithm: We developed the Quantum Entanglement Trees (QET) algorithm, leveraging the concept of local orderliness to optimize the QEM problem.
+  • Proposed two optimizations: We introduced the use of residuals to reduce MSE and employed batch processing techniques to accelerate QET algorithm.
 2 PROBLEM SETTING
@@
 Therefore, searching the entire space is impractical, and we need heuristic algorithms to prioritize exploring more promising regions of the search space.
+Figure 1: QET algorithm.
 3 METHODS
-The core idea of HETA (HETA) is to leverage the local orderliness of matrix elements by rearranging adjacent elements to optimize the matrix quantization algorithm.
+The core idea of Quantum Entanglement Trees (QET) is to leverage the local orderliness of matrix elements by rearranging adjacent elements to optimize the matrix quantization algorithm.
 Additionally, multiple iterations are used to expand the algorithm’s swapping field.
@@
 3.1 Basic algorithm
-The steps and design of the HETA algorithm are illustrated in Figure 1. The figure provides a conceptual overview of the HETA (HETA) algorithm.
-On the left side of the figure, the process of rearranging matrix elements through multiple iterations in the HETA is depicted.
+The steps and design of the QET algorithm are illustrated in Figure 2. The figure provides a conceptual overview of the Quantum Entanglement Trees (QET) algorithm.
+On the left side of the figure, the process of rearranging matrix elements through multiple iterations in the QET is depicted.
 This iterative procedure involves comparing adjacent elements and ensuring that adjacent elements are not placed in the same matrix.
-This separation of adjacent elements into different matrices is conceptually similar to quantum entanglement, hence the name ”HETA.” The HETA algorithm performs multiple iterations, with each iteration comparing adjacent elements and splitting them based on their values.
+This separation of adjacent elements into different matrices is conceptually similar to quantum entanglement, hence the name ”Quantum Entanglement Trees.” The QET algorithm performs multiple iterations, with each iteration comparing adjacent elements and splitting them based on their values.
 This approach is based on two observations.
@@
 This iterative process, similar to that of convolutional neural networks, extends the ”receptive field” to cover more elements.
-On the right side of the figure, the HETA compression process is shown.
+On the right side of the figure, the QET compression process is shown.
 After the elements have been rearranged, the algorithm groups the sub-matrices, followed by clustering to determine the centroids.
 These centroids are used to quantize the matrix.
-The final quantized matrix, codebook, and indicator maps are the outputs of the HETA algorithm.
+The final quantized matrix, codebook, and indicator maps are the outputs of the QET algorithm.
 This divide-and-conquer compression method is based on an observation: by splitting the matrix, the expressive power can be increased under a fixed storage constraint.
@@
 Without splitting, storing $c^{m}$ centroids would be required.
-3.2 HETA Quantization
-The proposed HETA Algorithm is designed to optimize the quantization of a matrix $X\in\mathbb{R}^{n\times d}$ by leveraging local orderliness.
+3.2 QET Quantization
+The proposed QET Algorithm is designed to optimize the quantization of a matrix $X\in\mathbb{R}^{n\times d}$ by leveraging local orderliness.
 The algorithm proceeds in four key steps, as shown in Algorithm 2:
@@
 The output of the algorithm includes the quantized matrix $X^{q}$, the codebook $\mathbf{C}$ containing the centroids, and the Indicator Maps $IM$.
-Algorithm 1 HETA Quantization Algorithm
-1:Input: Matrix $X\in\mathbb{R}^{n\times d}$, $m$ is the number of subspaces for quantization, $l$ is the number of HETA iterations.
+Algorithm 1 QET Quantization Algorithm
+1:Input: Matrix $X\in\mathbb{R}^{n\times d}$, $m$ is the number of subspaces for quantization, $l$ is the number of QET iterations.
 2:Output: Quantized matrix $X^{q}=\{X^{q}_{i,k}\}$ where $i=1,\dots,d$ and $k=1,\dots,m$, Codebook $\mathbf{C}=\{C_{i}\}$ where $i=1,\dots,m$, and Indicator Maps $IM=\{I_{i}\}$ where $i=1,\dots,l$
@@
 32:Return Quantized matrix $X^{q}$, Codebook $\mathbf{C}$, and Indicator Maps $IM$
-3.2.1 HETA Dequantization
-The dequantization process, as described by the HETA Dequantization Algorithm, aims to reconstruct the original matrix $X^{\prime}$ from its quantized version $X^{q}$.
+3.2.1 QET Dequantization
+The dequantization process, as described by the QET Dequantization Algorithm, aims to reconstruct the original matrix $X^{\prime}$ from its quantized version $X^{q}$.
 This process involves several key steps, as shown in Algorithm 2:
@@
 The final output of the algorithm is the reconstructed matrix $X^{\prime}\in\mathbb{R}^{n\times d}$, which approximates the original matrix before quantization.
-Algorithm 2 HETA Dequantization Algorithm
+Algorithm 2 QET Dequantization Algorithm
 1:Input: Quantized matrix $X^{q}=\{X^{q}_{i,k}\}$, where $i=1,\dots,d$ and $k=1,\dots,m$; Codebook $\mathbf{C}=\{C_{i}\}$, where $i=1,\dots,m$; and Indicator Maps $IM=\{I_{i}\}$, where $i=1,\dots,l$
@@
 In this section, we propose three optimizations: Residual Optimization (RO), Batch Processing Optimization (BPO), and Codebook Quantization Optimization (CQO).
-RO takes advantage of the fact that, after applying HETA, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE.
+RO takes advantage of the fact that, after applying QET, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE.
 On the other hand, BPO utilizes batch processing techniques to enhance processing speed.
-Since the codebook used in the HETA algorithm occupies significant space, CQO further quantizes the codebook to improve space efficiency.
+Since the codebook used in the QET algorithm occupies significant space, CQO further quantizes the codebook to improve space efficiency.
 Table 1: Range Ratios, and Codebook(CB) Proportions in different compression ratios.
@@
 3.3.1 Residual Optimization
-Residual Optimization (RO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the HETA algorithm.
+Residual Optimization (RO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the QET algorithm.
 As shown in Table 1, our experiments indicate that at a compression ratio of 16, the data range of the quantized matrix is reduced to 29.7% of the original matrix’s range.
@@
 The residual matrix typically exhibits a significantly reduced data range, making it easier to compress.
-This residual matrix is then quantized using a chosen matrix quantization algorithm, such as RTN or HETA.
+This residual matrix is then quantized using a chosen matrix quantization algorithm, such as RTN or QET.
 During the dequantization process, the quantized residual is dequantized, and this dequantized residual is added back to the initially dequantized matrix to produce a revised dequantized matrix.
@@
 3.3.2 Batch Processing Optimization
-The key insight behind (Batch Processing Optimization) BPO is that by organizing and processing data in larger chunks or batches and enhancing the efficiency of the HETA algorithm through batch processing and vectorization operations, the algorithm can more efficiently determine the relationships between elements, such as identifying smaller and larger parts or setting indicator maps.
+The key insight behind (Batch Processing Optimization) BPO is that by organizing and processing data in larger chunks or batches and enhancing the efficiency of the QET algorithm through batch processing and vectorization operations, the algorithm can more efficiently determine the relationships between elements, such as identifying smaller and larger parts or setting indicator maps.
 In the original process, elements of the matrix are compared and sorted individually, leading to a sequential, element-by-element approach that can become computationally intensive as the matrix size increases.
@@
 3.3.3 Codebook Quantization Optimization (CQO)
-The codebook occupies a significant portion of the space in the HETA algorithm.
+The codebook occupies a significant portion of the space in the QET algorithm.
 As shown in Table 1, our experiments indicate that when the compression ratio is 16, the codebook occupies 62.5% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space.
@@
 Our optimization mainly employs the RTN method to compress the codebook.
-4 Theory
+4 Theoretical Guarantees of QET
+In this section, we will first derive the MSE of the PQ and QET algorithms, then theoretically prove that the MSE of our QET algorithm is lower than that of the PQ algorithm.
+Theorem 1.
+Suppose each element in the matrix is independently sampled from a normal distribution $x_{(i,j)}$ $\sim$ $\mathcal{N}(\mu,\sigma)$. After Quantizated by the PQ algorithm, the MSE is as follows:
+$\text{MSE}=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$
+$A_{ij}\sim\mathcal{N}(\mu,\sigma^{2})\quad\text{independently and identically distributed}$
+[…]
+$C_{ij}\quad\text{residual matrix}$
+$E[C_{ij}]=E[A_{ij}-B_{ij}]=\mu-\mu=0$
+$\text{MSE}=\text{Var}[C_{ij}]=\text{Var}[A_{ij}]+\text{Var}[B_{ij}]-2\cdot\text{Cov}(A_{ij},B_{ij})$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\text{Cov}(A_{ij},B_{ij})$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(A_{ij}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}A_{ij}-\mu\right)\right]$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\left[\frac{1}{n_{k}}E\left[(A_{ij}-\mu)^{2}\right]+\frac{1}{n_{k}}\sum_{i\neq i^{\prime}}E\left[(A_{ij}-\mu)(A_{i^{\prime}j}-\mu)\right]\right]$
+$\text{Since}\ A_{ij}\text{ and}\ A_{i^{\prime}j}\ (i\neq i^{\prime})\ \text{are independent}$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}$
+$\text{MSE}=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$ (1)
+In this section, this paper will theoretically prove that the MSE of our HERA algorithm is lower than that of the PQ algorithm.
+Suppose each element in the matrix is independently sampled from a normal distribution $\mathcal{N}(\mu,\sigma)$. After compression by the PQ algorithm, the process for calculating MSE is as follows:
+$A_{ij}\sim\mathcal{N}(\mu,\sigma^{2})\quad\text{independently and identically distributed}$
+[…]
+$C_{ij}\quad\text{residual matrix}$
+$E[C_{ij}]=E[A_{ij}-B_{ij}]=\mu-\mu=0$
+$MSE=\text{Var}[C_{ij}]=\text{Var}[A_{ij}]+\text{Var}[B_{ij}]-2\cdot\text{Cov}(A_{ij},B_{ij})$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\text{Cov}(A_{ij},B_{ij})$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(A_{ij}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}A_{ij}-\mu\right)\right]$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\left[\frac{1}{n_{k}}E\left[(A_{ij}-\mu)^{2}\right]+\frac{1}{n_{k}}\sum_{i\neq i^{\prime}}E\left[(A_{ij}-\mu)(A_{i^{\prime}j}-\mu)\right]\right]$
+$\text{Since}\ A_{ij}\text{ and}\ A_{i^{\prime}j}\ (i\neq i^{\prime})\ \text{are independent}$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}$
+$=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$ (2)
+For the HERA algorithm proposed in this paper, the MSE obtained using the HERA algorithm on a matrix with elements following a $\mathcal{N}(\mu,\sigma^{2})$ distribution is equivalent to the MSE obtained using the PQ algorithm on a matrix with elements following a distribution with variance $0.682\sigma^{2}$.
+According to formula equation 2 equation 3, the MSE of HERA is smaller than that of the PQ .
+The following discussion in this paper will explain why this equivalence holds.
+For a Gaussian distribution, after applying element pairing in the HERA algorithm, the cumulative distribution function $F_{Y}(y)$ of the elements on the right side of the matrix (the side storing the larger elements) satisfies:
+$F_{Y}(y)=P(Y\leq y)=P(\max(X_{1},X_{2})\leq y)=P(X_{1}\leq y)\cdot P(X_{2}\leq y)$
+$\text{Each}\ X_{i}\text{ follows a normal distribution}\ N\ (\mu,\sigma^{2})$
+$F_{Y}(y)=\Phi\left(\frac{y-\mu}{\sigma}\right)\cdot\Phi\left(\frac{y-\mu}{\sigma}\right)=\left[\Phi\left(\frac{y-\mu}{\sigma}\right)\right]^{2}=\frac{1}{4}\left[1+\text{erf}\left(\frac{y-\mu}{\sigma\sqrt{2}}\right)\right]^{2}$
+Take the derivative of $\text{Take the derivative of }F_{Y}(y):$
+$f_{Y}(y)=\left[1+\text{erf}\left(\frac{y-\mu}{\sigma\sqrt{2}}\right)\right]\cdot\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)$
+The mean and variance of $\text{The mean and variance of }Y\text{ are given by:}$:
+[…]
+[…]
+Use numerical integration methods:
+$E(Y)=\mu+0.564\sigma$
+$Var(Y)=0.682\sigma^{2}$ (3)
+Taking symmetry into account, the left side of the matrix also adheres to equation equation 3.
 5 Experiments
@@
 Due to space constraints, we present the results for the synthetic normal distribution dataset and the LLM weight dataset in the main text.
-The results for the other datasets are provided in the appendix (Section A.1).
+The results for the other datasets are provided in the appendix (Section LABEL:sec:appexponaddi).
 Below, we provide a detailed description of each dataset.
@@
 (4) Gist dataset: The Gist dataset consists of feature vectors used for image recognition and retrieval tasks.
-This dataset is particularly useful for evaluating the algorithm’s effectiveness in handling high-dimensional and dense data representations.
-(Jegou et al.
-(2010)) (Oliva & Torralba (2001))
-(5) Wiki dataset: The Wiki dataset contains document embeddings derived from Wikipedia articles. This dataset tests the algorithm’s ability to compress and process large-scale text data efficiently.
+GIST consists of the first 100k features extracted from the tiny image set (Torralba et al.
+(2008)).
+The dataset is also used in Jegou et al.
+(2010) and Oliva & Torralba (2001).
+(5) Wiki dataset: The Wiki dataset (wik (2023)) contains document embeddings derived from Wikipedia articles. This dataset includes 256-dimensional vectors for a sample of 1 million Wikipedia entries.
 Platform and implementation: We conducted our algorithm evaluations on a high-performance server equipped with an Intel Core i9-10980XE processor, featuring 18 cores and 36 threads, operating at a base frequency of 3.00 GHz.
@@
 In each repetition, matrices of the same size and from the same distribution were generated using different seeds.
-Metrics:
-Compared algorithms:
-Comparison of Algorithms and Parameter Selection
+Metrics: We primarily measure the accuracy and time consumption of the algorithm.
+We use MAE (Mean Absolute Error), MRE (Mean Relative Error), and MSE (Mean Squared Error) as accuracy metrics.
+QTR (Quantization Time Ratio) and DTR (Dequantization Time Ratio) are used as time consumption metrics.
+Below, we introduce these metrics in detail.
+Let $x_{(i,j)}$ denote the elements of the original matrix to be quantized, and $x_{(i,j)}^{\prime}$ denote the elements of the dequantized matrix.
+$\text{MAE}=\frac{1}{n\cdot d}\sum_{i,j}\left|x_{(i,j)}-x_{(i,j)}^{\prime}\right|,\;\text{MSE}=\frac{1}{n\cdot d}\sum_{i,j}\left(x_{(i,j)}-x_{(i,j)}^{\prime}\right)^{2}$
+For time efficiency, we define the following metrics:
+$\text{QTR}=\frac{T_{\text{quant}}}{T_{\text{quant, baseline}}},\;\text{DTR}=\frac{T_{\text{dequant}}}{T_{\text{dequant, baseline}}}$
+In these equations, $T_{\text{quant}}$ and $T_{\text{dequant}}$ represent the time taken by the proposed algorithm to quantize and dequantize the data, respectively.
+$T_{\text{quant, baseline}}$ and $T_{\text{dequant, baseline}}$ refer to the corresponding times taken by the baseline algorithm.
+Comparative Algorithms: For the abstract Quantization Error Minimization (QEM) problem, our comparative algorithms fall into two main categories.
+The first category involves independently compressing the elements of a matrix for each specific scenario, which can be abstracted as the RTN (Round-To-Nearest) algorithm.
+The second category groups matrix elements by columns and then applies quantization to each group.
+Related algorithms in this category include PQ, OPQ, and LOPQ.
+Therefore, our comparative algorithms are as follows: RTN (Round-To-Nearest) (Gray & Neuhoff (1998)), PQ (Product Quantization) (Jegou et al.
+(2010)), OPQ (Optimized Product Quantization) (Ge et al.
+(2013)), and LOPQ (Locally Optimized Product Quantization) (Kalantidis & Avrithis (2014)).
+Parameter Selection
 In this section, we describe the parameter selection process for our HERA algorithm and compare its performance with the Optimized Product Quantization (OPQ) algorithm.
-Our HERA algorithm involves a critical parameter, the number of centroids $k_{s}$. The selection of $k_{s}$ is governed by the following memory constraint:
-$k_{s}\times d\times 32+n\times m\times[(k_{s}-1)\cdot\text{bitlength}]\leq\text{memory in bits}$
+For Residual Optimization (RO), the large space required by RTN makes it infeasible for the algorithm to achieve high compression ratios.
+Therefore, in our experiments, we primarily use QET as the residual algorithm for RO.
+Our QET algorithm involves a critical parameter, the number of centroids $k_{s}$. The selection of $k_{s}$ is governed by the following memory constraint:
 where:
@@
 The OPQ algorithm optimizes product quantization by adjusting the space partitioning and centroid assignments to minimize quantization error.
-Comparison Metrics
-We use MAE (Mean Absolute Error), MRE (Mean Relative Error), and MSE (Mean Squared Error) for experimental evaluation.
-Let $x_{(i,j)}$ and $x_{(i,j)}^{\prime}$ denote the values before and after dequantization, respectively.
-The following metrics are used for comparison:
-$\text{MAE}=\frac{1}{n\cdot d}\sum_{i,j}\left|x_{(i,j)}-x_{(i,j)}^{\prime}\right|$
-$\text{MRE}=\frac{1}{n\cdot d}\sum_{i,j}\frac{\left|x_{(i,j)}-x_{(i,j)}^{\prime}\right|}{x_{(i,j)}}$
-$\text{MSE}=\frac{1}{n\cdot d}\sum_{i,j}\left(X_{(i,j)}-X_{(i,j)}^{\prime}\right)^{2}$
-We performed an optimization using 1-4 iterations, corresponding to our algorithms labeled as our1 through our4.
-In the quantification accuracy experiment, we tested the accuracy with different numbers of groups $m$ and various sizes of data matrices.
-In the sensitivity experiment, we evaluated the impact of different numbers of groups $m$ on the results while maintaining the same compression ratio.
 Accuracy Measurement
@@
 Figure 26: MSE on 4 iterations.
-We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  4-  4, Figures  7-  7, Figures  10-  10, Figures  13-  13 respectively.
+We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  5-  5, Figures  8-  8, Figures  11-  11, Figures  14-  14 respectively.
 For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
@@
 We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance.
-The results for 1 to 4 iterations are shown in Figures  1616, Figures  1919, and Figures  $2$, respectively.
+The results for 1 to 4 iterations are shown in Figures  1717, Figures  2020, and Figures  2323, respectively.
 Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).
+6 Future Work
+In future work, we plan to apply this algorithm to the quantization of large model parameters and the kv cache.
+This will allow us to evaluate the algorithm’s effectiveness in handling the challenges posed by large-scale models and real-time processing scenarios, further extending its applicability and performance in practical machine learning tasks.
 7 Conclusion
@@
 The promising results of this study encourage continued research and development in the quest for more effective and scalable solutions for LLM compression and optimization.
-Appendix A Appendix
-A.1 Experimental Results on Additional Datasets

### v4

@@
-Quantum Entanglement Trees: Optimizing Quantized Matrix Quantization via Element Replacement and Residual Clustering
+QET: Enhancing Quantized LLM Parameters and KV cache Compression through Element Substitution and Residual Clustering
 Abstract
@@
 To enhance QET, we introduce two optimizations: further quantizing residuals to reduce MSE, and using masking and batch processing to accelerate the algorithm.
-Experimental results demonstrate that QET can effectively reduce MSE to 12.3% of the original value at the same compression ratio, outperforming the best baseline algorithms.
+Experimental results demonstrate that QET can effectively reduce MSE to 5.05%, 13.33%, and 11.89% of the current best method on the LLM dataset, K cache, and V cache, respectively.
 Our contributions include the abstraction of the QEM problem, the design of the QET algorithm, and the proposal of two optimizations to improve accuracy and speed.
@@
 We evaluate QET on multiple datasets.
-Experimental results demonstrate that our method can effectively reduce the MSE to 12.3% of the original value at the same compression ratio, compared to the best baseline algorithm, which achieves an MSE of XXXX%.
+Experimental results demonstrate that QET can effectively reduce MSE to 5.05%, 13.33%, and 11.89% of the current best method on the LLM dataset, K cache, and V cache, respectively.
 We summarize our contributions below.
@@
   • Designed an algorithm: We developed the Quantum Entanglement Trees (QET) algorithm, leveraging the concept of local orderliness to optimize the QEM problem.
-  • Proposed two optimizations: We introduced the use of residuals to reduce MSE and employed batch processing techniques to accelerate QET algorithm.
+  • Proposed two optimizations: We introduced residual quantization to reduce MSE and codebook quantization techniques to optimize memory efficiency.
 2 PROBLEM SETTING
@@
 18:Return Reconstructed matrix $X^{\prime}$
+3.3 Theoretical Guarantees of QET
+In this section, we will first derive the MSE of the PQ and QET algorithms, then theoretically prove that the MSE of our QET algorithm is lower than that of the PQ algorithm.
+Theorem 1.
+Suppose each element in the matrix is independently sampled with a mean $\mu$ and variance $\sigma^{2}$. After quantization by the PQ algorithm, the Mean Squared Error (MSE) is given by:
+$\text{MSE}_{\text{PQ}}=\sigma^{2}\left(1-\frac{1}{n_{k}}\right).$
+Proof.
+Let $c_{(i,j)}$ denote the cluster centroid matrix, and let $n_{k}$ be the number of vectors in the cluster. Define the residual matrix as $s_{(i,j)}$.
+The expectation of the residual matrix is:
+$E[s_{(i,j)}]=E[x_{(i,j)}-c_{(i,j)}]=\mu-\mu=0.$
+The Mean Squared Error (MSE) can be expressed as the variance of the residual matrix:
+$\text{MSE}_{\text{PQ}}=\text{Var}[s_{(i,j)}]=\text{Var}[x_{(i,j)}]+\text{Var}[c_{(i,j)}]-2\cdot\text{Cov}(x_{(i,j)},c_{(i,j)}).$
+Given that $c_{(i,j)}$ is the centroid within the cluster:
+$\text{MSE}_{\text{PQ}}$ $=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(x_{(i,j)}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}x_{(i,j)}-\mu\right)\right]$
+$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{1}{n_{k}}\left(E\left[(x_{(i,j)}-\mu)^{2}\right]+\sum_{i\neq i^{\prime}}E\left[(x_{(i,j)}-\mu)(x_{i^{\prime}j}-\mu)\right]\right).$
+Since $x_{(i,j)}$ and $x_{i^{\prime}j}$ (for $i\neq i^{\prime}$) are independent:
+$\text{MSE}_{\text{PQ}}=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}=\left(1-\frac{1}{n_{k}}\right)\sigma^{2}.$ (1)
+Theorem 2.
+For the QET algorithm without optimization, where matrix elements are independently sampled from a normal distribution $x_{(i,j)}\sim\mathcal{N}(\mu,\sigma^{2})$, the Mean Squared Error (MSE) is:
+$\text{MSE}_{\text{QET}}=0.682\times\text{MSE}_{\text{PQ}}.$
+Proof.
+We denote by $F_{Y}(y)$ the cumulative distribution function (CDF) of the elements on the right side of the matrix (i.e., the larger elements).
+$F_{Y}(y)=P(Y\leq y)=P(\max(X_{l},X_{r})\leq y)=P(X_{l}\leq y)\cdot P(X_{r}\leq y)$
+Where $y$ represents the elements of the matrix formed by the larger of the adjacent elements in matrix $X$.
+$X_{l}$ and $X_{r}$ denote the random variables corresponding to the adjacent left and right elements in matrix $X$, respectively.
+Since $X_{l}$ and $X_{r}$ follow a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, we have:
+$F_{Y}(y)=\Phi\left(\frac{y-\mu}{\sigma}\right)\cdot\Phi\left(\frac{y-\mu}{\sigma}\right)=\left[\Phi\left(\frac{y-\mu}{\sigma}\right)\right]^{2}=\frac{1}{4}\left[1+\text{erf}\left(\frac{y-\mu}{\sigma\sqrt{2}}\right)\right]^{2}$
+Take the derivative of $F_{Y}(y)$ with respect to $y$:
+$f_{Y}(y)=\left[1+\text{erf}\left(\frac{y-\mu}{\sigma\sqrt{2}}\right)\right]\cdot\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)$
+The mean and variance of $Y$ are given by:
+[…]
+[…]
+Using numerical integration methods, we obtain:
+$\mathbb{E}(Y)=\mu+0.564\sigma,\quad\text{Var}(Y)=0.682\sigma^{2}.$
+Because QET can be considered a special case of the PQ algorithm, according to Equation 1,
+$\text{MSE}_{\text{QET}}=\left(1-\frac{1}{n_{k}}\right)\text{Var}(Y)=0.682\left(1-\frac{1}{n_{k}}\right)\sigma^{2}=0.682\cdot\text{MSE}_{\text{PQ}}.$ (2)
+Taking symmetry into account, this conclusion also holds for the smaller matrix on the left. ∎
+Theorem 1 demonstrates that the MSE of quantized matrix is proportional to the variance of the distribution.
+For any distribution, as long as the variance of the matrix elements is reduced, the MSE will also decrease.
+Theorem 2 shows that when the matrix elements follow a normal distribution, the MSE is reduced to 0.682 times its original value.
 3.4 Optimizations
-In this section, we propose three optimizations: Residual Optimization (RO), Batch Processing Optimization (BPO), and Codebook Quantization Optimization (CQO).
-RO takes advantage of the fact that, after applying QET, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE.
-On the other hand, BPO utilizes batch processing techniques to enhance processing speed.
-Since the codebook used in the QET algorithm occupies significant space, CQO further quantizes the codebook to improve space efficiency.
+In this section, we propose three optimizations: Residual Quantization Optimization (RQO) and Codebook Quantization Optimization (CQO).
+RQO takes advantage of the fact that, after applying QET, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE.
+CQO further quantizes the codebook to improve space efficiency.
 Table 1: Range Ratios, and Codebook(CB) Proportions in different compression ratios.
-Compression Ratio 4.0 8.0 12.0 16.0
-Range Ratio (%) 1.6 4.0 9.2 29.7
-CB Proportion (%) 87.5 78.1 71.8 62.5
-3.3.1 Residual Optimization
-Residual Optimization (RO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the QET algorithm.
-As shown in Table 1, our experiments indicate that at a compression ratio of 16, the data range of the quantized matrix is reduced to 29.7% of the original matrix’s range.
+Compression Ratio 4.0 8.0 12.0
+Range Ratio (%) 1.6 4.0 9.2
+CB Proportion (%) 87.5 78.1 71.8
+3.4.1 Residual Quantization Optimization
+Residual Quantization Optimization (RQO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the QET algorithm.
+As shown in Table 1, our experiments indicate that at a compression ratio [^3] Compression ratio = Space occupied by the matrix before quantizationSpace occupied by the matrix after $\frac{\text{Space occupied by the matrix before quantization}}{\text{Space occupied by the matrix after quantization}}$ of 12, the data range of the quantized matrix is reduced to 9.2% of the original matrix’s range.
 Moreover, with a compression ratio of 4, the data range of the quantized matrix is reduced to approximately 1.6% of the original matrix’s range.
-Matrix quantization algorithms are more efficient at compressing matrix with a smaller data range; therefore, we propose Residual Optimization (RO).
-The RO algorithm, as depicted in Figure 2, begins with the quantization of the original matrix to produce a quantized matrix.
+Matrix quantization algorithms are more efficient at compressing matrix with a smaller data range; therefore, we propose Residual Quantization Optimization (RQO).
+The RQO algorithm, as depicted in Figure 2, begins with the quantization of the original matrix to produce a quantized matrix.
 This matrix is then dequantized to yield a dequantized matrix, which is subtracted from the original matrix to form a residual matrix.
@@
 Figure 2: Residual Optimization Algorithm.
-3.3.2 Batch Processing Optimization
-The key insight behind (Batch Processing Optimization) BPO is that by organizing and processing data in larger chunks or batches and enhancing the efficiency of the QET algorithm through batch processing and vectorization operations, the algorithm can more efficiently determine the relationships between elements, such as identifying smaller and larger parts or setting indicator maps.
-In the original process, elements of the matrix are compared and sorted individually, leading to a sequential, element-by-element approach that can become computationally intensive as the matrix size increases.
-This method is not optimal for large-scale matrices.
-BPO introduces a more advanced technique by processing matrix elements in parallel batches rather than sequentially.
-This parallelization allows the algorithm to handle multiple comparisons and sorting operations simultaneously, rather than one at a time.
-By doing so, it reduces the overall computational complexity and accelerates the entire quantization process.
-3.3.3 Codebook Quantization Optimization (CQO)
+3.4.2 Codebook Quantization Optimization
 The codebook occupies a significant portion of the space in the QET algorithm.
-As shown in Table 1, our experiments indicate that when the compression ratio is 16, the codebook occupies 62.5% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space.
+As shown in Table 1, our experiments indicate that when the compression ratio is 12, the codebook occupies 71.8% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space.
 Therefore, compressing the codebook can significantly reduce the space used by the algorithm.
+We proposed Codebook Quantization Optimization (CQO).
 Our optimization mainly employs the RTN method to compress the codebook.
-4 Theoretical Guarantees of QET
-In this section, we will first derive the MSE of the PQ and QET algorithms, then theoretically prove that the MSE of our QET algorithm is lower than that of the PQ algorithm.
-Theorem 1.
-Suppose each element in the matrix is independently sampled from a normal distribution $x_{(i,j)}$ $\sim$ $\mathcal{N}(\mu,\sigma)$. After Quantizated by the PQ algorithm, the MSE is as follows:
-$\text{MSE}=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$
-$A_{ij}\sim\mathcal{N}(\mu,\sigma^{2})\quad\text{independently and identically distributed}$
-[…]
-$C_{ij}\quad\text{residual matrix}$
-$E[C_{ij}]=E[A_{ij}-B_{ij}]=\mu-\mu=0$
-$\text{MSE}=\text{Var}[C_{ij}]=\text{Var}[A_{ij}]+\text{Var}[B_{ij}]-2\cdot\text{Cov}(A_{ij},B_{ij})$
-$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\text{Cov}(A_{ij},B_{ij})$
-$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(A_{ij}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}A_{ij}-\mu\right)\right]$
-$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\left[\frac{1}{n_{k}}E\left[(A_{ij}-\mu)^{2}\right]+\frac{1}{n_{k}}\sum_{i\neq i^{\prime}}E\left[(A_{ij}-\mu)(A_{i^{\prime}j}-\mu)\right]\right]$
-$\text{Since}\ A_{ij}\text{ and}\ A_{i^{\prime}j}\ (i\neq i^{\prime})\ \text{are independent}$
-$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}$
-$\text{MSE}=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$ (1)
-In this section, this paper will theoretically prove that the MSE of our HERA algorithm is lower than that of the PQ algorithm.
-Suppose each element in the matrix is independently sampled from a normal distribution $\mathcal{N}(\mu,\sigma)$. After compression by the PQ algorithm, the process for calculating MSE is as follows:
-$A_{ij}\sim\mathcal{N}(\mu,\sigma^{2})\quad\text{independently and identically distributed}$
-[…]
-$C_{ij}\quad\text{residual matrix}$
-$E[C_{ij}]=E[A_{ij}-B_{ij}]=\mu-\mu=0$
-$MSE=\text{Var}[C_{ij}]=\text{Var}[A_{ij}]+\text{Var}[B_{ij}]-2\cdot\text{Cov}(A_{ij},B_{ij})$
-$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\text{Cov}(A_{ij},B_{ij})$
-$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(A_{ij}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}A_{ij}-\mu\right)\right]$
-$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\left[\frac{1}{n_{k}}E\left[(A_{ij}-\mu)^{2}\right]+\frac{1}{n_{k}}\sum_{i\neq i^{\prime}}E\left[(A_{ij}-\mu)(A_{i^{\prime}j}-\mu)\right]\right]$
-$\text{Since}\ A_{ij}\text{ and}\ A_{i^{\prime}j}\ (i\neq i^{\prime})\ \text{are independent}$
-$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}$
-$=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$ (2)
-For the HERA algorithm proposed in this paper, the MSE obtained using the HERA algorithm on a matrix with elements following a $\mathcal{N}(\mu,\sigma^{2})$ distribution is equivalent to the MSE obtained using the PQ algorithm on a matrix with elements following a distribution with variance $0.682\sigma^{2}$.
-According to formula equation 2 equation 3, the MSE of HERA is smaller than that of the PQ .
-The following discussion in this paper will explain why this equivalence holds.
-For a Gaussian distribution, after applying element pairing in the HERA algorithm, the cumulative distribution function $F_{Y}(y)$ of the elements on the right side of the matrix (the side storing the larger elements) satisfies:
-$F_{Y}(y)=P(Y\leq y)=P(\max(X_{1},X_{2})\leq y)=P(X_{1}\leq y)\cdot P(X_{2}\leq y)$
-$\text{Each}\ X_{i}\text{ follows a normal distribution}\ N\ (\mu,\sigma^{2})$
-$F_{Y}(y)=\Phi\left(\frac{y-\mu}{\sigma}\right)\cdot\Phi\left(\frac{y-\mu}{\sigma}\right)=\left[\Phi\left(\frac{y-\mu}{\sigma}\right)\right]^{2}=\frac{1}{4}\left[1+\text{erf}\left(\frac{y-\mu}{\sigma\sqrt{2}}\right)\right]^{2}$
-Take the derivative of $\text{Take the derivative of }F_{Y}(y):$
-$f_{Y}(y)=\left[1+\text{erf}\left(\frac{y-\mu}{\sigma\sqrt{2}}\right)\right]\cdot\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)$
-The mean and variance of $\text{The mean and variance of }Y\text{ are given by:}$:
-[…]
-[…]
-Use numerical integration methods:
-$E(Y)=\mu+0.564\sigma$
-$Var(Y)=0.682\sigma^{2}$ (3)
-Taking symmetry into account, the left side of the matrix also adheres to equation equation 3.
 4 Experiments
@@
 Below, we provide a detailed description of each dataset.
+Each element in the matrix is represented with 32 bits, except for the KV cache dataset, where each element is represented with 16 bits.
 (1) Synthetic normal distribution dataset: This dataset was generated by drawing each element of the matrix from a truncated normal distribution with a mean of 0.5 and a standard deviation of 0.16.
 Additionally, one out of every ten thousand elements was replaced with an outlier value, randomly chosen between -100 and 100.
-(2) LLM weight dataset: This dataset comprises weight matrices extracted from the large language model (LLM) LLaMA2 (Touvron et al. (2023)).
+For the synthetic normal distribution dataset, we generated three types of matrices of different sizes: synthetic dataset 1: 1024 $\times$ 128 matrices, synthetic dataset 2: 1024 $\times$ 512 matrices, and synthetic dataset 3: 1024 $\times$ 1024 matrices.
+(2) LLM weight dataset: This dataset comprises weight matrices extracted from the large language model (LLM) LLaMA2 (Touvron et al.
+(2023)).
+The sizes of the LLM weight matrices are 11008 $\times$ 4096.
 (3) KV cache dataset: The KV cache dataset is derived from the key-value pairs stored in the cache during inference in KV Quant (Hooper et al.
@@
 The key and value are stored in matrices, respectively.
-(4) Gist dataset: The Gist dataset consists of feature vectors used for image recognition and retrieval tasks.
-GIST consists of the first 100k features extracted from the tiny image set (Torralba et al.
-(2008)).
-The dataset is also used in Jegou et al.
-(2010) and Oliva & Torralba (2001).
-(5) Wiki dataset: The Wiki dataset (wik (2023)) contains document embeddings derived from Wikipedia articles. This dataset includes 256-dimensional vectors for a sample of 1 million Wikipedia entries.
-Platform and implementation: We conducted our algorithm evaluations on a high-performance server equipped with an Intel Core i9-10980XE processor, featuring 18 cores and 36 threads, operating at a base frequency of 3.00 GHz.
-The server also includes 128GB of 3200MHz DDR4 memory and a 24.8MB L3 cache, providing robust computational capabilities.
-All algorithms were implemented in Python, using version 3.8.10.
-For each case, the experiment was repeated 100 times.
-In each repetition, matrices of the same size and from the same distribution were generated using different seeds.
+In the KV cache dataset, the sizes of both the K matrices and the V matrices are 4096 $\times$ 4096.
+Platform and implementation: We conducted our algorithm evaluations on a high-performance server equipped with an Intel Xeon Platinum 6462C (Sapphire Rapids) processor, featuring 16 virtual CPUs (vCPUs), operating at a base frequency of 3.3 GHz, with a maximum turbo frequency of 3.9 GHz.
+The server also includes 64GB of memory, providing robust computational capabilities.
+All algorithms were implemented and executed on this server environment to ensure optimal performance for our experimental evaluations.
 Metrics: We primarily measure the accuracy and time consumption of the algorithm.
 We use MAE (Mean Absolute Error), MRE (Mean Relative Error), and MSE (Mean Squared Error) as accuracy metrics.
-QTR (Quantization Time Ratio) and DTR (Dequantization Time Ratio) are used as time consumption metrics.
+Quantization time ratio and Dequantization time ratio are used as time consumption metrics.
 Below, we introduce these metrics in detail.
@@
 $\text{MAE}=\frac{1}{n\cdot d}\sum_{i,j}\left|x_{(i,j)}-x_{(i,j)}^{\prime}\right|,\;\text{MSE}=\frac{1}{n\cdot d}\sum_{i,j}\left(x_{(i,j)}-x_{(i,j)}^{\prime}\right)^{2}$
-For time efficiency, we define the following metrics:
-$\text{QTR}=\frac{T_{\text{quant}}}{T_{\text{quant, baseline}}},\;\text{DTR}=\frac{T_{\text{dequant}}}{T_{\text{dequant, baseline}}}$
-In these equations, $T_{\text{quant}}$ and $T_{\text{dequant}}$ represent the time taken by the proposed algorithm to quantize and dequantize the data, respectively.
-$T_{\text{quant, baseline}}$ and $T_{\text{dequant, baseline}}$ refer to the corresponding times taken by the baseline algorithm.
+For time efficiency, we use Quantization Time (QT) and DeQuantization Time (DQT) as the metrics.
 Comparative Algorithms: For the abstract Quantization Error Minimization (QEM) problem, our comparative algorithms fall into two main categories.
@@
 (2013)), and LOPQ (Locally Optimized Product Quantization) (Kalantidis & Avrithis (2014)).
-Parameter Selection
-In this section, we describe the parameter selection process for our HERA algorithm and compare its performance with the Optimized Product Quantization (OPQ) algorithm.
-For Residual Optimization (RO), the large space required by RTN makes it infeasible for the algorithm to achieve high compression ratios.
-Therefore, in our experiments, we primarily use QET as the residual algorithm for RO.
-Our QET algorithm involves a critical parameter, the number of centroids $k_{s}$. The selection of $k_{s}$ is governed by the following memory constraint:
-where:
-  • $k_{s}$ is the number of centroids,
-  • $d$ is the dimensionality of each sample,
-  • $n$ is the number of samples,
-  • $m$ is the number of subspaces,
-  • bitlength is the bit length used for encoding.
-This constraint ensures that the total memory usage remains within the available memory limits.
-By carefully selecting $k_{s}$, we aim to optimize the performance of the HERA algorithm while adhering to memory constraints.
-The Optimized Product Quantization (OPQ) algorithm was used as a benchmark for comparison.
-The OPQ algorithm was implemented with default parameters, providing a consistent basis for performance evaluation.
-The OPQ algorithm optimizes product quantization by adjusting the space partitioning and centroid assignments to minimize quantization error.
-Accuracy Measurement
-Figure 3: MAE on 8 subspaces.
-Figure 4: MRE on 8 subspaces.
-Figure 5: MSE on 8 subspaces.
-Figure 6: MAE on 16 subspaces.
-Figure 7: MRE on 16 subspaces.
-Figure 8: MSE on 16 subspaces.
-Figure 9: MAE on 32 subspaces.
-Figure 10: MRE on 32 subspaces.
-Figure 11: MSE on 32 subspaces.
-Figure 12: MAE on 128 subspaces.
-Figure 13: MRE on 128 subspaces.
-Figure 14: MSE on 128 subspaces.
-Figure 15: MAE on 1 iterations.
-Figure 16: MRE on 1 iterations.
-Figure 17: MSE on 1 iterations.
-Figure 18: MAE on 2 iterations.
-Figure 19: MRE on 2 iterations.
-Figure 20: MSE on 2 iterations.
-Figure 21: MAE on 3 iterations.
-Figure 22: MRE on 3 iterations.
-Figure 23: MSE on 3 iterations.
-Figure 24: MAE on 4 iterations.
-Figure 25: MRE on 4 iterations.
-Figure 26: MSE on 4 iterations.
-We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  5-  5, Figures  8-  8, Figures  11-  11, Figures  14-  14 respectively.
-For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
-From the experimental results, we can derive four conclusions.
-Firstly, the higher the compression ratio, the less space is consumed, but the accuracy decreases.
-Secondly, the OPQ algorithm performs worse than others with the same compression rate because it needs to store the transformation matrix.
-Additionally, the PQ algorithm does not perform as well as our improved algorithm.
-Finally, the more iterations, the better the performance of our algorithm.
-We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance.
-The results for 1 to 4 iterations are shown in Figures  1717, Figures  2020, and Figures  2323, respectively.
-Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).
+Parameter Selection: We first introduce the parameter settings of the QET algorithm.
+For an $n\times d$ matrix with $a$ bit per element, we perform a quantization operation with a compression ratio of $\theta$.
+The QET algorithm iterates $l$ rounds.
+For residual optimization, we perform $r$ residual iterations.
+During the $i$th residual step, the number of clusters for centroids is $k_{i}$, and the number of groups during the grouping operation is $m$.
+For codebook quantization optimization, the bit length of the codebook after compression is $a^{\prime}$ bit.
+$R_{i}$ denotes the ratio of the space occupied by the codebook and the quantized matrix to the total space excluding the indicator maps during the $i$th iteration of residual optimization.
+To ensure that the total memory usage remains within the available memory limits, we have the following constraints on the parameters of the QET algorithm.
+$k_{i}\times d\times a^{\prime}+2a+n\times m\times\log_{2}(k_{i})\leq\left[\frac{n\times d\times a}{\theta}-0.5\times n\times d\times l\right]\times R_{i},\quad\text{for }i=1,\ldots,r.$
+Given the parameters $d$, $a^{\prime}$, $a$, $n$, $m$, $\theta$, $l$, and $R_{i}$, we calculate the value of $k_{i}$ using the bisection method.
+Unless otherwise specified, the QET parameters are configured as follows: $n$, $d$, and $a$ are determined by the input matrix.
+We perform a standard QET and one residual QET, both with codebook quantization.
+$r=2$, the first QET has 3 iterations, and the second QET does not perform matrix permutation ($l=3$).
+Additionally, $d/m=8$, $a^{\prime}=10$, $R_{1}=70\%$, and $R_{2}=30\%$.
+For RTN, PQ, OPQ, and LOPQ, under the same space constraints, we use the same common parameters as our algorithm, while other parameters are set according to the recommended configurations from their respective papers.
+We define the names of different versions of the algorithm as follows: Vanilla refers to the basic QET algorithm without any optimizations.
+Vanilla_CQO refers to the Vanilla algorithm with codebook quantization optimization, where the codebook is quantized using the RTN algorithm.
+Vanilla_RQO indicates the Vanilla algorithm with residual optimization, where the residuals are processed using the QET algorithm.
+Finally, QET is the Vanilla algorithm with both residual optimization and codebook quantization optimization (including all optimizations), where the codebook is quantized using the RTN algorithm and the residuals are processed using the QET algorithm.
+4.2 Matrix quantization result
+In this section, we provide a comprehensive comparison of the accuracy and computational efficiency of the proposed QET algorithm against other benchmark methods.
+We will first discuss the precision metrics (MAE and MSE) across different datasets and compression ratios, followed by an analysis of the time metrics (QT and DQT) to evaluate the efficiency of each algorithm.
+The goal is to demonstrate the effectiveness of the QET algorithm in balancing accuracy with computational speed across various real and synthetic datasets.
+Some algorithms are unable to run on certain datasets or at specific compression ratios.
+For instance, in the case of square matrices, the rotation matrices for OPQ and LOPQ become as large as the original matrix, making compression impossible.
+In such cases, we observe missing data points or curves in the corresponding figures where the algorithms could not produce results.
+(a) MAE of Synthetic 1.
+(b) MSE of Synthetic 1.
+(c) MAE of Synthetic 2.
+(d) MSE of Synthetic 2.
+(e) MAE of Synthetic 3.
+(f) MSE of Synthetic 3.
+(g) MAE of LLM.
+(h) MSE of LLM.
+(i) MAE of K cache.
+(j) MSE of K cache.
+(k) MAE of V cache.
+(l) MSE of V cache.
+Figure 3: MAE and MSE of different datasets.
+MAE and MSE In our experiments, we evaluated the performance of our quantization algorithm across multiple datasets using MAE and MSE.
+These metrics provide a comprehensive understanding of the algorithm’s accuracy at different compression ratios, ranging from 2 to 16.
+The average MSE of RTN for the synthetic dataset 1, synthetic dataset 2, synthetic dataset 3, LLM dataset, K cache dataset, and V cache dataset is 71.74 times, 334.59 times, 247.39 times, 12.56 times, 10.57 times, and 7.69 times that of QET, respectively.
+Therefore, we do not include RTN in the figures for comparison.
+For synthetic dataset 1: As shown in Figures  3(a) and  3(b) , our QET algorithm consistently outperforms the baseline algorithms across all compression ratios.
+The MAE and MSE values increase as the compression ratio rises, but QET maintains a significantly lower error compared to PQ and OPQ algorithms.
+The Vanilla version, a faster variant of QET, shows slightly higher error than the full QET version, but still remains competitive.
+For example, at a compression ratio of 4, compared to the best-performing algorithm, PQ, QET reduces the MSE to 6.94%, while the Vanilla version reduces the MSE to 36.53%.
+For Synthetic Dataset 2: In Figures  3(c) and  3(d), a similar trend is observed with Synthetic Dataset 2.
+The QET algorithm achieves the lowest MAE and MSE values, demonstrating its robustness across different types of synthetic data.
+Notably, the gap between QET and the other methods becomes more pronounced as the compression ratio increases, indicating the superiority of QET in handling higher compression rates without significantly sacrificing accuracy.
+Due to the space consumption of OPQ and LOPQ, they were unable to handle the compression of square matrices in this dataset.
+The vanilla version performs similarly to the QET algorithm.
+For example, at a compression ratio of 4, compared to the best-performing algorithm, PQ, QET reduces the MSE to 7.13%, while the Vanilla version reduces the MSE to 36.51%.
+For Synthetic Dataset 3: Figures  3(e) and  3(f) show the results for the third synthetic dataset.
+While the errors increase with higher compression ratios, QET still outperforms the other algorithms.
+The difference in performance becomes especially significant at higher compression ratios (e.g., 12 to 16), where QET and vanilla continue to show lower error rates.
+For example, at a compression ratio of 4, compared to the best-performing algorithm, PQ, QET reduces the MSE to 7.21%, while the Vanilla version reduces the MSE to 36.54%.
+For the LLM dataset, as depicted in Figures  3(g) and  3(h), the QET algorithm demonstrates superior performance.
+The MAE and MSE values for QET are consistently lower than those for Vanilla, PQ, and OPQ, indicating that QET is well-suited for compressing large-scale language model weights without significantly sacrificing accuracy.
+At higher compression ratios (13 to 16), the MAE and MSE of Vanilla are similar to those of QET.
+For example, at a compression ratio of 4, compared to the best-performing algorithm, PQ, QET reduces the MSE to 5.05%, while the Vanilla version reduces the MSE to 36.64%.
+For KV Cache Dataset: The results of K cache is shown in Figures  3(i) and  3(j).
+The results of V cache is shown in Figures  3(k) and  3(l).
+Similar to the LLM datasets, the QET algorithm achieves the best results, with lower MAE and MSE values across all compression ratios.
+At compression ratios between 13 and 16, the MAE and MSE values of Vanilla closely approach those of QET.
+For example, at a compression ratio of 4 for K cache, compared to the best-performing algorithm, PQ, QET reduces the MSE to 13.33%, while the Vanilla version reduces the MSE to 50.24%.
+And at a compression ratio of 4 for V cache, compared to the best-performing algorithm, PQ, QET reduces the MSE to 11.89%, while the Vanilla version reduces the MSE to 45.83%.
+In summary, the QET algorithm consistently outperforms baseline methods, achieving lower MAE and MSE values across all datasets and compression ratios.
+Vanilla also demonstrates relatively low MAE and MSE values, and on certain datasets, it performs comparably to QET at higher compression ratios.
+(a) QT of Synthetic 1.
+(b) DQT of Synthetic 1.
+(c) QT of Synthetic 2.
+(d) DQT of Synthetic 2.
+(e) QT of Synthetic 3.
+(f) DQT of Synthetic 3.
+(g) QT of LLM.
+(h) DQT of LLM.
+(i) QT of K cache.
+(j) DQT of K cache.
+(k) QT of V cache.
+(l) DQT of V cache.
+Figure 4: QT (s) and DQT (s) of different datasets.
+QT and DQT In this experiments, we evaluated the performance of the quantization process by measuring both Quantization Time (QT) and Dequantization Time (DQT) across multiple datasets, as shown in Figures 4.
+These metrics provide a comprehensive understanding of the computational efficiency of the algorithm at different compression ratios, ranging from 2 to 16.
+For QT, the QET algorithm exhibits longer quantization times at lower compression ratios, but as the compression ratio increases, QT decreases significantly.
+This phenomenon is expected, as fewer centroids simplify the clustering process, resulting in faster compression as the compression ratio increases.
+Vanilla consistently achieves lower QT across all datasets compared to QET, making it the faster option, albeit with a slight loss in accuracy.
+In contrast, PQ and other baseline methods demonstrate moderate QT performance, although OPQ and LOPQ show slower performance at larger matrix sizes due to their complexity.
+As for DQT, QET maintains a relatively stable performance across all compression ratios and datasets, with marginal increases in time as the compression ratio grows.
+Vanilla follows a similar trend but achieves faster dequantization.
+PQ remains competitive in terms of both QT and DQT.
+4.3 Ablation Studies and Discussions
+These experiments demonstrate the effectiveness of the algorithm optimizations.
+Specifically, both Residual Quantization Optimization (RQO) and Codebook Quantization Optimization (CQO) improve the baseline Vanilla algorithm and provide comparisons with the QET algorithm.
+The ablation studies primarily focus on accuracy improvements.
+We evaluated the performance on the LLM dataset by measuring both MAE and MSE.
+(a) MAE of RQO.
+(b) MSE of RQO.
+(c) MAE of CQO.
+(d) MSE of CQO.
+Figure 5: Ablation Studies of RQO and CQO.
+Residual Quantization Optimization (RQO) As illustrated in Figures 5(b) and LABEL:fig:ablationrqomse, the introduction of RQO into the Vanilla algorithm leads to a notable reduction in MAE and MSE values.
+At lower compression ratios, Vanilla with RQO achieves performance that is very close to QET, particularly in terms of MSE.
+The gap widens at higher compression ratios because, while the storage space remains the same for both algorithms, RQO requires more codebooks due to the multi-layer structure, which reduces the number of centroids in clustering.
+On the other hand, QET optimizes codebook compression, allowing for more centroids and better MAE and MSE performance at higher compression ratios.
+For instance, at a compression ratio of 4, the RQO optimization reduces MSE to 16.23% of the original value.
+Codebook Quantization Optimization (CQO)
+In Figures 5(c) and 5(d), the impact of CQO on the Vanilla algorithm is shown.
+CQO helps to decrease MAE and MSE, especially at lower compression ratios.
+CQO optimization effectively reduces the size of the codebook, leading to improvements in both MAE and MSE.
+However, at higher compression ratios, the accuracy improvements become less noticeable because the codebook size is already small at these higher compression levels, leaving limited room for further compression.
+For instance, at a compression ratio of 4, the CQO optimization reduces MSE to 40.08% of the original value.
 5 Future Work
@@
 6 Conclusion
-In conclusion, our research addresses critical challenges in the optimization of large language models (LLMs) through the development of a novel algorithm, HERA.
-The growing complexity and scale of LLMs necessitate innovative approaches to enhance storage and computational efficiency, particularly in vector database management, weight quantization, and key-value (k-v) cache optimization.
-HERA’s hierarchical approach to segmenting, compressing, and reorganizing the matrix dataset has proven to be effective in significantly reducing quantization error.
-By considering the distribution and magnitude of parameters, our method achieves superior performance compared to traditional uniform quantization techniques.
-The experimental results from our prototype system, implemented in Python, demonstrate that HERA can reduce the quantization error to 12.3% of the original, maintaining the same compression ratio.
-Future work will explore further refinements to HERA, including its application to other types of application and broader datasets.
-Additionally, investigating the integration of HERA with other optimization techniques may yield even greater enhancements in model performance and efficiency.
-The promising results of this study encourage continued research and development in the quest for more effective and scalable solutions for LLM compression and optimization.
+Matrix quantization involves representing matrix elements in a more space-efficient form to reduce storage usage, with dequantization restoring the original matrix for practical applications.
+We frame the Quantization Error Minimization (QEM) problem as minimizing the distance between a matrix before and after quantization, constrained by the condition that the quantized matrix occupies the same memory space.
+Matrix quantization plays a crucial role in various domains, including Large Language Models (LLMs) weight quantization, vector databases, KV cache quantization, graph compression, and image compression.
+Recent advancements in LLMs, such as GPT-4 and BERT, have underscored the importance of matrix compression due to the large size of parameters and KV caches, which are stored as matrices.
+To address the QEM problem, we propose Quantum Entanglement Trees (QET), which leverage the local orderliness of matrix elements through iterative element swapping to form a locally ordered matrix.
+This matrix is subsequently grouped and quantized by columns.
+To enhance the QET algorithm, we introduce two key optimizations: further quantizing residuals to reduce Mean Squared Error (MSE) and utilizing masking and batch processing techniques to accelerate the algorithm.
+Our experimental results demonstrate that QET can effectively reduce MSE to 5.05%, 13.33%, and 11.89% of the best existing methods on the LLM dataset, K cache, and V cache, respectively.
+These findings highlight QET’s efficiency in reducing quantization error while maintaining compression effectiveness.
+The contributions of this work include the formal abstraction of the QEM problem, the design of the QET algorithm, and the introduction of two optimizations that improve both accuracy and computational speed.
+Future work will focus on extending this algorithm to the quantization of large-scale model parameters and KV caches.
+This will allow us to further evaluate the algorithm’s effectiveness in addressing the challenges posed by large-scale models and real-time processing scenarios, thereby broadening its applicability and impact in practical machine learning tasks.